### v5

# Foundation Models on a Budget: Approximating Blocks in Large Vision Models

Abstract
Foundation Models have shown impressive performance in various tasks and domains, yet they require massive computational resources, raising concerns about accessibility and sustainability. Previous attempts to reduce foundation model size fall short of fully addressing the problem, as they end up increasing computational load through additional training steps. Recent works reveal that deep neural networks exhibit internal representation similarities. While inter-network similarities have enabled techniques such as model stitching and merging, intra-network similarities remain underexplored for improving efficiency. In this paper, we propose Transformer Blocks Approximation (TBA), a novel method that leverages intra-network similarities to identify and approximate transformer blocks in large vision models. TBA replaces these blocks using lightweight, closed-form transformations, without retraining or fine-tuning the rest of the model. The proposed method reduces the number of parameters while having minimal impact on the downstream task. We validate the effectiveness and generalizability of TBA through extensive experiments across multiple datasets (e.g., Imagenet-1k and CIFAR100) and state-of-the-art pretrained vision models (e.g, ViT, DiNO-v2, and DEiT).
1 Introduction
As Neural Networks (NNs) grow in size and complexity, their demand for computational resources has become a significant bottleneck. Despite the impressive performance of large models, they often come with substantial trade-offs, such as increased memory and power consumption. This has led to a growing interest in methods that can reduce model complexity without sacrificing performance. However, most existing approaches trying to mitigate these challenges either demand additional training steps or result in substantial performance degradation. Recent research showed that there exist similarities between representation within and between NNs. The observation that many layers or components within these networks perform similar functions or produce highly correlated outputs suggests that some blocks may be approximated, highlighting opportunities for network simplification. In this paper, we investigate two key research questions: (i) how to effectively identify blocks within neural networks that yield similar representations, and (ii) how to approximate these blocks with minimal impact on overall performance without any additional training steps. To address the first question, we provide an extensive analysis showing that the Mean Squared Error (MSE) effectively identifies such blocks, suggesting that their approximation should have a negligible impact. For the second question, we propose Transformer Blocks Approximation (TBA), a novel method that leverages similarities between representations to approximate transformer blocks using lightweight transformations, such as linear mappings. Once the blocks that have minimal impact on model functionality are identified, TBA replaces them with a simpler transformation. This approximation strategy allows for reducing the overall parameter count, ensuring minimal impact on downstream task performance. Our main contributions are as follows:
We introduce TBA, a general framework for identifying and approximating transformer blocks that produce similar outputs, using simpler transformations (e.g., linear), reducing model parameters and complexity with minimal to no impact on the downstream task performance (Figure 1).
2 Related work
Measuring similarities A range of metrics have been introduced to assess the similarity between latent spaces generated by different NNs (Klabunde et al., 2023; Ballester et al., 2023). One established approach is Canonical Correlation Analysis (CCA) (Hotelling, 1992), known for its invariance to linear transformations. Variants of CCA, such as Singular Value CCA (SVCCA) (Raghu et al., 2017), aim to enhance robustness, while techniques like Projection Weighted CCA (PWCCA) (Morcos et al., 2018) mitigate sensitivity to small perturbations. Another widely used metric, Centered Kernel Alignment (CKA) (Kornblith et al., 2019), captures the similarity between latent spaces while ignoring orthogonal transformations. However, recent work (Davari et al., 2022) highlights that this metric can be sensitive to shifts in the latent space. Additionally, Barannikov et al. (2021) proposes a method to compare two data representations by measuring the multi-scale topological dissimilarity, while Fumero et al. (2024) leverages the principles of spectral geometry to model and analyze the relationships between distinct latent spaces.
Leveraging similarities Valeriani et al. (2024) examine the intrinsic dimensionality and neighbor compositions of representations in transformer models. Kvinge et al. (2022) investigate how models process variations in data points across layers, while Nguyen et al. (2020) assess the impact of network depth and width on hidden representations. Additionally, Crisostomi et al. (2023) study the conditions under which two latent spaces can be merged into a unified one. Moschella et al. (2023) construct a unified space shared by different NNs, enabling zero-shot stitching of independently trained models across different modalities (Norelli et al., 2023). More recently, Cannistraci et al. (2024) enable model stitching without explicit assumptions about the transformation class connecting the latent manifold embeddings, or with only partial correspondence between latent spaces (Cannistraci et al., 2023). Finally, Lähner & Moeller (2024); Maiorca et al. (2024) demonstrate that representations learned by distinct NNs can be aligned using simple transformations.
Architectural efficiency While large-scale models with billions or even trillions of parameters continue to achieve state-of-the-art performance, their growth comes with trade-offs, such as slower inference times and significantly higher computational costs. Improving the efficiency of Deep Neural Network (DNN) has been a long-standing area of research. For instance, Veit et al. (2016) shows that removing residual blocks from deep Convolutional Neural Networks (CNNs) only marginally impacts performance, which inspired approaches to reduce inference time by dynamically deciding which layers to execute based on the input (Wu et al., 2018; Veit & Belongie, 2018). Additionally, various techniques to enhance efficiency have emerged, such as early exiting and model pruning. Early exit strategies, which introduce intermediate output layers at different stages of the network, have been shown to reduce inference time (Xin et al., 2020; Zhou et al., 2020; Yu et al., 2022; Tang et al., 2023). However, these approaches require the training of intermediate classifiers to enable exits at predefined layers. Alternatively, model pruning reduces computational load by either removing individual weights based on specific criteria, such as gradient information (Ma et al., 2023), entropy (Liao et al., 2023), or second-order information (Singh & Alistarh, 2020), or by eliminating larger structural components, like channels or residual blocks in ResNets (Bai et al., 2023; Wang & Wu, 2023) and self-attention layers in Transformers (Zhang & He, 2020; Sajjad et al., 2023; Venkataramanan et al., 2024; Zhang et al., 2024). Although effective, these approaches require training the model from scratch and, in the best case, fine-tuning. However, Bai et al. (2023) shows that for CNNs, this additional training step can sometimes be avoided.
Unlike other methods, TBA leverages intra-network similarities to reduce foundation model complexity without the need for additional training steps while maintaining competitive performance.
3 Transformer Blocks Approximation
The core idea of the proposed approach is to detect similar representations within NNs and replace the corresponding blocks with simple transformations. In the context of this work, a "block" refers to a sequence of layers including multi-head self-attention, normalization, and feed-forward layers, functioning as a cohesive unit. Instead of executing the entire NN, using transformer block approximation (TBA), we approximate these blocks, reducing computational complexity while preserving the network’s functionality. To provide a clear understanding, Figure 1 presents an overview of our method.
Identifying similar representations We hypothesize that certain foundation models may contain blocks that produce similar representations. To quantify these similarities, we employ the MSE. A low MSE between the output of an earlier block and a later block indicates that their respective representations are highly similar. This suggests that the intervening sequence of blocks contributes minimally to transforming the representation, highlighting potential redundancy.
Let represent the total number of blocks in the model. Let denote the output representation (a vector of dimension ) of block for a single input, where . For any two blocks, a starting block and an ending block (where ), we compute their output representations and . For each , where is a subset of points sampled uniformly at random from , defines the input to the corresponding layer . The MSE between these representations is defined as:
By systematically evaluating the MSE for various pairs of blocks , we can identify those being good candidates for approximation. This allows for targeted complexity reduction without substantially altering the network’s internal representations or downstream-task performance at a very low computational cost.
Approximating transformer blocks Once two blocks and are identified as having highly similar output representations, our goal is to replace the intermediate blocks with a single, lightweight transformation that maps the output of block directly to an approximation of the output of block . This approach allows us to skip the computation of blocks , effectively reducing the overall computational costs. This approximation can be repeated for multiple, non-overlapping blocks, i.e., blocks and with .
Let and represent the output representations from block and respectively, for the data points in . Our objective is to find a transformation such that:
In this work, we consider to be a linear transformation that can be estimated by minimizing the squared error between the transformed output and the actual , which can be solved using least squares:
This optimization problem allows for a closed-form solution that efficiently computes the optimal transformation . The solution bypasses the computation of all layers between any two blocks and , replacing them with . This approximation significantly reduces computational complexity by replacing one or more transformer blocks, comprising multi-head self-attention and feed-forward layers, with a low-cost linear transformation. To sum up, our approach consists of two main stages:
1.
Identification: We use MSE to identify blocks that produce similar representations.
2.
Approximation: We compute an efficient linear approximation using a transformation matrix to bypass the previously identified blocks.
This procedure reduces parameter count with minimal impact on internal representations, as shown in Figures 9, 7, 10, 11 and 8, and maintains compatibility with any downstream classifiers, achieving significant compression as shown Tables 1, 8, 9, 10, 11, 12 and 13.
4 Experiments
This section begins with an analysis of block-level similarities within pre-trained vision foundation models (Section 4.1), and we investigate how TBA impacts the representations of these latent spaces. Following this, Section 4.2 provides a comprehensive quantitative evaluation of TBA’s performance on various image classification tasks, including standard, zero-shot, and cross-dataset generalization scenarios. Lastly, in Section 4.3 we compare TBA against existing baselines, including an ablation study on transformation complexity, further validating the effectiveness of the proposed approach.
4.1 Latent analysis
Block similarities We investigate similarities emerging in latent representations using three distinct transformer-based models with different dimensionality: ViT-S, DiNO-B, and DEiT-S, and five datasets: CIFAR-10, CIFAR-100, MNIST, F-MNIST, and ImageNet1k.
Details on the models and datasets are provided in Table 6 and Table 7. Figure 2 presents the MSE matrices between blocks of different models when using different datasets. These similarities are calculated using only a small subset of the training data (i.e., 500 samples). Results reveal that while the patterns of similarity vary across models, they remain consistent across different datasets. This suggests that the similarity structure between computational blocks is predominantly influenced by the model rather than the dataset used. This finding aligns with observations from Nguyen et al. (2020), where DNNs trained from scratch tend to exhibit a distinctive “block structure” in their representations, linked to model overparameterization. Our results extend this observation to pretrained foundation models, where we show that such a structure is primarily influenced by the model. We conduct the analysis using the mean over the tokens, but in Section A.2.1 we show that this also applies when using only the CLS token from each block. Finally, in Figure 6, we present an ablation study to investigate the influence of the similarity metric. We correlate various metrics with downstream task performance using ViT-S and ImageNet1k. Our findings reveal that using simple metrics such as MSE or cosine similarity to identify block-level similarities exhibits a strong correlation with final accuracy. These results underscore the flexibility of our proposed method, indicating its compatibility with a range of similarity or distance functions that can capture inter-block relationships.
Effects of TBA on latent representations After identifying blocks producing similar representations, we analyze the impact on the latent representation output by the last block when approximating using TBA.
These blocks are approximated using a shared linear transformation applied across all tokens based on a subset of 3,000 training samples. For consistency, we employed the same models and datasets used for the results in Figure 2. We compute the MSE between the representations in the last block of the original and the TBA model when approximating the block. As illustrated in Figure 3, in most cases, the MSE increases as the block depth increases.
Original TBA
DiNO-B
DEiT-S
This suggests that approximating the final blocks would lead to significant changes in the final representations, indicating their critical role. However, in the case of DEiT-S, the trend is different. Here, the MSE decreases toward the final blocks. This is also shown in the similarity matrices illustrated in Figure 2, where the final blocks have a high similarity. These findings reinforce the intuition behind the use of MSE, demonstrating a correlation between the metric and the final representation similarity when approximating blocks. In some instances, such as with the MNIST dataset, the MSE scores remain relatively consistent across blocks, indicating that the representations are largely similar to one another. However, for more complex datasets like ImageNet1k, the representations in the final or the first blocks become increasingly dissimilar, making it advantageous to approximate intermediate blocks. This suggests that the similarities are primarily driven by the model but also guided by the complexity of the dataset, allowing for targeted approximations that reduce model parameters and complexity without significantly compromising performance. To further investigate the effect of TBA on the latent representations, we visualize, using Principal Component Analysis (PCA), the representations of the TBA model when approximating the final block and its original representation (see Figure 4). The plot shows the representations generated using the DiNO-B and DEiT-S pretrained encoders on F-MNIST. Colors represent classes. The visualization highlights that approximating the final block of DiNO-B results in noticeable deviations from the original representations, while in DEiT-S the approximated representations remain similar to the original ones. This observation aligns with the results from Figure 3, where approximating blocks can lead to significant changes in representations. For complete results, refer to Section A.2.2.
4.2 Transformer block approximation performance
We perform image classification using various pretrained models with different dimensionalities (ViT-T, ViT-S, ViT-L, ViT-B, DiNO-S, DiNO-B, and DEiT-S) and datasets (MNIST, F-MNIST, CIFAR-10, CIFAR-100C, CIFAR-100F, and ImageNet1k), keeping all models frozen. Block approximations are computed using a shared linear transformation applied across all tokens, based on
| Accuracy | |||||
| Approx. | Params. | CIFAR-10 | CIFAR-100F | ImageNet1k | |
| ViT-S | 1 5 | -6.51M | (-11.09) | (-27.35) | (-39.86) |
| 2 5 | -4.88M | (-4.88) | (-14.42) | (-17.56) | |
| 7 10 | -4.88M | (-8.69) | (-23.89) | (-51.14) | |
| 2 4 | -3.26M | (-2.95) | (-8.67) | (-6.80) | |
| 9 11 | -3.26M | (-5.00) | (-14.71) | (-35.54) | |
| 1 2, 4 5 | -3.26M | (-3.54) | (-11.10) | (-12.06) | |
| 0 1 | -1.63M | (-2.09) | (-6.15) | (-11.10) | |
| 3 4 | -1.63M | (-0.66) | (-2.47) | (-2.51) | |
| 9 10 | -1.63M | (-1.15) | (-6.19) | (-15.58) | |
| 10 11 | -1.63M | (-1.67) | (-4.97) | (-12.73) | |
| original | 21.82M | ||||
| DiNO-B | 1 5 | -26.5M | (-11.28) | (-28.62) | (-58.57) |
| 2 5 | -19.5M | (-5.87) | (-19.50) | (-28.26) | |
| 7 10 | -19.5M | (-4.46) | (-14.02) | (-35.88) | |
| 2 4 | -13M | (-1.49) | (-6.62) | (-7.70) | |
| 9 11 | -13M | (-10.76) | (-14.36) | (-27.74) | |
| 1 2, 4 5 | -13M | (-1.47) | (-6.11) | (-7.79) | |
| 0 1 | -6.5M | (-0.05) | (-1.01) | (-1.00) | |
| 3 4 | -6.5M | (-0.38) | (-2.78) | (-1.62) | |
| 9 10 | -6.5M | (-0.78) | (-3.26) | (-10.09) | |
| 10 11 | -6.5M | (-1.01) | (-4.23) | (-13.83) | |
| original | 86.58M | ||||
| DEiT-S | 1 5 | -6.51M | (-13.55) | (-28.90) | (-40.88) |
| 2 5 | -4.88M | (-5.41) | (-14.87) | (-16.05) | |
| 7 10 | -4.88M | (-1.61) | (-2.78) | (-22.24) | |
| 2 4 | -3.26M | (-1.85) | (-6.37) | (-5.32) | |
| 9 11 | -3.26M | (+0.30) | (+1.11) | (-5.39) | |
| 1 2, 4 5 | -3.26M | (-5.74) | (-13.31) | (-10.68) | |
| 0 1 | -1.63M | (-6.21) | (-12.91) | (-15.38) | |
| 3 4 | -1.63M | (-0.14) | (-1.24) | (-1.22) | |
| 9 10 | -1.63M | (+0.30) | (+0.86) | (-2.15) | |
| 10 11 | -1.63M | (+0.49) | (+1.15) | (+0.05) | |
| original | 21.82M | ||||
a subset of 3,000 training samples. Then, we train a single linear layer using the Adam optimizer (learning rate ) for 5 epochs, with 3 seeds and a batch size of 256. Details regarding the models, datasets, computational resources, and software tools are provided in Table 6, Table 7, Section A.1.4, and Section A.1.3, respectively.
Image classification As presented in Table 1, the proposed TBA method successfully reduces model size while maintaining, and in some cases even improving, image classification performance. These downstream task results support the findings from the analysis in Section 4.1. Specifically, as discussed in Section 4.1 and illustrated in Figures 3 and 4, approximating the final block of DEiT-S yields a final representation highly similar to the original. This similarity suggests such blocks are optimal candidates for approximation, a conclusion further validated by Table 1. The table shows that even when approximating multiple consecutive blocks (e.g., 911), we can reduce parameters while achieving performance superior to the original model. Overall, the consistent or enhanced performance indicates that a linear transformation is sufficient to approximate these transformer blocks, thereby effectively reducing model parameters. It’s important to note that this transformation is shared across all tokens, with no additional training steps required. Additional results are presented in Tables 8, 9, 10, 11, 12 and 13. To assess the scalability of our method, we extended our evaluation to ViT-L. While our approach reduces its parameters, ViT-L still imposes considerable computational demands, necessitating high-performance GPU infrastructure. This finding underscores that while effective for substantial compression, our method’s impact on making extremely large models practical for resource-constrained environments is limited by their initial scale. Finally, in Section A.2.6, we further analyze model behavior after block approximations on the final image-classification task, complementing the quantitative results with qualitative insights into the approximation procedure.
Zero-shot image classification To further assess the effectiveness of our approach, we evaluate
| Params. | Approx. | Accuracy | |
| -6.49M | 0 1 | 57.93 | - |
| 1 2 | 64.20 | - | |
| 2 3 | 66.35 | -5.51% | |
| 3 4 | 64.65 | - | |
| 4 5 | 64.86 | - | |
| 5 6 | 58.05 | - | |
| 6 7 | 61.56 | - | |
| 7 8 | 58.53 | - | |
| 8 9 | 52.32 | - | |
| 9 10 | 59.21 | - | |
| 10 11 | 22.64 | - | |
| 149.07M | original | 70.21 | – |
TBA in a zero-shot image classification setting. This evaluation utilizes the OpenCLIP-ViT-B model (Radford et al., 2021), which was pretrained on LAION-2B Schuhmann et al. (2022), with ImageNet1k serving as the downstream evaluation dataset. As in previous experiments, the model remains frozen, and block approximations are computed using a shared linear transformation applied across all tokens, based on a subset of 3,000 training samples. Importantly, we apply these approximations only to the vision encoder, leaving the text encoder unchanged. We follow the standard ImageNet1k prompt templates. As shown in Table 2, when approximating earlier vision blocks (e.g., 1 2 or 2 3), the model still achieves competitive results, but slightly reduces the number of parameters. The analysis is conducted only on the base version, as larger versions (e.g., OpenCLIP-ViT-L or OpenCLIP-ViT-H) contain too many parameters and are thus beyond the scope of this paper. Details on the models are provided in Table 6.
Transformation generalization We further evaluate the model’s ability to generalize transformations across datasets. Specifically, we analyze the capability of applying a linear transformation learned on one dataset (source) to another (target), using the same underlying pretrained architecture and without any subsequent training or fine-tuning of the approximated model. Table 3 reports accuracy scores for ViT-S and DiNO-S on MNIST, CIFAR-10, and CIFAR-100, where the "Source" column indicates the source dataset used to compute the transformation. The results demonstrate that a single, token-shared linear transformation can generalize effectively across different target datasets, with the exception of MNIST, which may be too simplistic to yield broadly transferable representations. This cross-dataset generalization is particularly valuable in scenarios where fine-tuning is infeasible, such as in privacy-sensitive domains like healthcare or finance. In these contexts, regulatory constraints often restrict training on local data, yet leveraging existing pretrained models remains crucial. Our findings suggest that lightweight adaptation via shared linear transformations offers a promising path for improving model performance on custom datasets without requiring additional training. Additional results are provided in Tables 14 and 15.
| Accuracy | ||||
| Approx. | Source | MNIST | CIFAR-10 | CIFAR-100C |
| 3 4 | MNIST | |||
| CIFAR-10 | ||||
| CIFAR-100F | ||||
| 3 5 | MNIST | |||
| CIFAR-10 | ||||
| CIFAR-100F | ||||
4.3 Comparison
Direct comparisons for TBA are limited, as existing approaches either fine-tune pretrained models or train from scratch. In contrast, TBA maintains the foundation model frozen. We include two types of comparisons: a variant of SkipAt, a method proposed in Venkataramanan et al. (2024), and an ablation study on transformation complexity.
SkipAt identity Venkataramanan et al. (2024) explores the effect of skipping self-attention or multi-head self-attention by inserting identity functions in pretrained models, without retraining. We extend this idea by skipping entire blocks, not just attention components. As shown in Table 4, TBA consistently outperforms this SkipAt Identity baseline on CIFAR-10 and CIFAR-100F. The accuracy improvements confirm that approximating is more effective than skipping them. Results with other datasets and models can be found in Section A.2.5.
| CIFAR-10 | CIFAR-100F | |||
| Approx. | SkipAt | TBA | SkipAt | TBA |
| 1 5 | ||||
| 2 5 | ||||
| 7 10 | ||||
| 1 3 | ||||
| 2 4 | ||||
| 9 11 | ||||
| 2 3 | ||||
| 9 10 | ||||
| original | ||||
Trained approximators We also evaluate whether more complex approximations improve performance. Specifically, we compare TBA to deeper MultiLayer Perceptron (MLP)-based translators (MLP and Res-MLP), which are trained for 300 steps using Adam with a learning rate of 1e-3. As shown in Table 5, employing a simple linear transformation to approximate blocks is the optimal choice in most cases. While deeper translators occasionally yield slightly higher accuracy (e.g., 15), they incur significant training overhead and require gradient-based optimization. In contrast, TBA is entirely training-free and operates in closed form; no backpropagation or updates are required. This highlights the efficiency–performance trade-off and supports the design choice behind TBA. Refer to Section A.1.2 for approximator details.
| Approx. | TBA | MLP | Res-MLP |
| 1 5 | |||
| 2 5 | |||
| 7 10 | |||
| 1 3 | |||
| 2 4 | |||
| 9 11 | |||
| 3 4 | |||
| 9 10 |
5 Limitations and future work
While TBA demonstrates significant promise in efficiently approximating transformer blocks, our current investigation has primarily focused on vision transformer architectures and their application to classification tasks. Future research will explore the applicability of TBA to other modalities (e.g., text) and to diverse downstream tasks (e.g., image reconstruction). Such an expansion will be crucial for testing the universality of the observed block-similarity phenomena and assessing TBA’s adaptability. Furthermore, we aim to expand the analysis of these block-level similarities. This involves investigating redundancies at finer granularities, such as within individual attention heads or feed-forward network layers, and potentially at coarser granularities across larger segments of the network. Another important direction is to examine more in detail how these similarity patterns arise depending on token type (e.g., CLS token versus last token). The analysis could lead to even more refined and context-aware approximation strategies, further enhancing model efficiency.
6 Conclusion
In this work, we first analyze the emergence of consistent block-wise representation similarities within pretrained foundation models and then propose a method to leverage these similarities to obtain smaller yet performant models. To this end, we propose Transformer Blocks Approximation (TBA), a novel method for identifying and efficiently approximating similar transformer blocks using a simple linear transformation, without requiring additional training or fine-tuning. Our extensive empirical evaluations across multiple pretrained vision models and datasets validate that TBA significantly reduces model parameters while maintaining, and sometimes even improving, downstream task performance. Notably, the learned linear transformations exhibit strong generalization capabilities across datasets, enabling lightweight model adaptation. Furthermore, TBA’s straightforward linear approach often achieves better results than existing strategies like block skipping, and can be more effective than complex, trained approximations. TBA thus offers a practical and efficient method for streamlining foundation models, making them more computationally accessible.
Acknowledgments
The authors gratefully acknowledge Luca Moschella for the insightful discussions. TS and EP are supported by the grant #2021-911 of the Strategic Focal Area “Personalized Health and Related Technologies (PHRT)” of the ETH Domain (Swiss Federal Institutes of Technology). EP is also supported by a fellowship from the ETH AI Center. This work has received funding from the Swiss State Secretariat for Education, Research, and Innovation (SERI).
Appendix A Appendix
A.1 Implementation details
This section details the experiments conducted in Section 4, providing information to reproduce them. Additionally, we provide the code as a zip file in the supplementary material.
A.1.1 Models and Datasets
Table 6 contains the full list of the pretrained models, while Table 7 contains dataset information.
| Modality | HuggingFace Model Name | Alias | Enc. Dim |
| Vision | WinKawaks/vit-tiny-patch16-224 | ViT-T (Dosovitskiy et al., 2021) | 192 |
| WinKawaks/vit-small-patch16-224 | ViT-S (Dosovitskiy et al., 2021) | 384 | |
| facebook/dinov2-small | DiNO-S (Oquab et al., 2023) | 384 | |
| facebook/deit-small-patch16-224 | DEiT-S (Touvron et al., 2020) | 384 | |
| google/vit-base-patch16-224 | ViT-B (Dosovitskiy et al., 2021) | 768 | |
| facebook/dinov2-base | DiNO-B (Oquab et al., 2023) | 768 | |
| laion/CLIP-ViT-B-16-laion2B-s34B-b88K | OpenCLIP-ViT-B (Zhai et al., 2019) | 768 | |
| google/vit-large-patch16-224 | ViT-L (Dosovitskiy et al., 2021) | 1024 |
| Modality | Name | Alias | Classes |
| Vision | MNIST (Deng, 2012) | MNIST | 10 |
| Fashion-MNIST (Xiao et al., 2017) | F-MNIST | 10 | |
| CIFAR-10 (Krizhevsky et al., 2009) | CIFAR-10 | 10 | |
| CIFAR-100 (coarse) (Krizhevsky et al., 2009) | CIFAR-100C | 20 | |
| CIFAR-100 (fine) (Krizhevsky et al., 2009) | CIFAR-100F | 100 | |
| Imagenet-1k (Russakovsky et al., 2015) | ImageNet1k | 1000 |
A.1.2 Approximators
The first implementation, referred to as the Res-MLP, is composed of two normalization layers and a feedforward submodule. The first layer normalization processes the input tensor, followed by a feedforward submodule comprising a linear transformation, a SiLU activation, a dropout layer, and a final linear transformation. The output of the feedforward submodule is added to the normalized input via a residual connection. This sum is then passed through the second normalization layer to produce the final output. The second implementation, referred to as the MLP, is a simplified MLP that employs a sequential architecture with a first linear transformation that reduces the input dimensionality to half of the target dimension, followed by a GELU activation function for smooth non-linearity, and a final linear transformation that restores the reduced features to match the target dimensionality. Refer to LABEL:lst:mlp1 and LABEL:lst:mlp2 for the code snippet of the two translators.
A.1.3 Tools & technologies
All the experiments presented in this work employ the following tools:
PyTorch Lightning, to ensure reproducible results while also getting a clean and modular codebase;
NN-Template GrokAI (2021), to easily bootstrap the project and enforce best practices;
Transformers by HuggingFace, to get ready-to-use transformers for both text and images;
Datasets by HuggingFace, to access most of the datasets;
DVC (Kuprieiev et al., 2022), for data versioning;
A.1.4 Computational resources
Experiments involving larger models, specifically DiNO-B, OpenCLIP-ViT-B, and ViT-L, were conducted on an NVIDIA H100 GPU equipped with 93 GB of memory. All the other experiments utilized an NVIDIA GeForce RTX 5090 GPU with 31 GB of memory.
A.2 Additional Experiments
This section presents supplementary experiments to extend those detailed in Section 4.
A.2.1 Block Similarities
The analysis in Section 4.1 reveals distinct block-wise similarity patterns within pretrained foundation models. The analysis utilizes the mean of token representations within each block to compute similarities. However, as detailed in Figure 5, these characteristic similarity patterns persist and are qualitatively very similar when using only the CLS token’s representation from each block. This suggests a degree of interchangeability for this type of analysis.
Finally, to assess the impact of the similarity measure itself, we conducted an ablation study correlating various metrics with downstream task performance on ViT-S and ImageNet1k (Figure 6). Our findings reveal that even simple metrics, such as MSE or cosine similarity, when used to identify block-level similarities, exhibit a strong correlation with final accuracy. These results underscore the robustness and flexibility of leveraging such block similarities, indicating compatibility with a range of similarity or distance functions capable of capturing these inter-block relationships.
| Approx. | Params. | ImageNet1k |
| 1 5 | 15.31M | |
| 2 5 | 16.94M | |
| 7 10 | 16.94M | |
| 1 3 | 18.56M | |
| 3 5 | 18.56M | |
| 2 4 | 18.56M | |
| 9 11 | 18.56M | |
| 2 3 | 20.19M | |
| 3 4 | 20.19M | |
| 4 5 | 20.19M | |
| 9 10 | 20.19M | |
| - | 21.82M |
A.2.2 PCA Visualization
In Section 4.1 we show the effect of TBA on latent representations. In this section, we provide additional visualization using PCA for DiNO-S, DEiT-S, ViT-S, with different datasets and approximating both early and late blocks (see Figures 10, 11, 9, 7 and 8).
A.2.3 Image Classification
This section presents additional experiments that complement and extend those detailed in Section 4.2. Datasets and models are the ones detailed in Tables 7 and 6.
| Approx. | Params. | MNIST | F-MNIST | CIFAR-10 | CIFAR-100C | CIFAR-100F | ImageNet1k |
| 1 5 | 15.31M | ||||||
| 2 5 | 16.94M | ||||||
| 7 10 | 16.94M | ||||||
| 1 3 | 18.56M | ||||||
| 2 4 | 18.56M | ||||||
| 3 5 | 18.56M | ||||||
| 1 2, 3 4 | 18.56M | ||||||
| 1 2, 4 5 | 18.56M | ||||||
| 0 1 | 20.43M | ||||||
| 1 2 | 20.43M | ||||||
| 2 3 | 20.43M | ||||||
| 3 4 | 20.43M | ||||||
| 4 5 | 20.43M | ||||||
| 5 6 | 20.43M | ||||||
| 6 7 | 20.43M | ||||||
| 7 8 | 20.43M | ||||||
| 8 9 | 20.43M | ||||||
| 9 10 | 20.43M | ||||||
| 10 11 | 20.43M | ||||||
| original | 22.06M |
| Approx. | Params. | MNIST | F-MNIST | CIFAR-10 | CIFAR-100C | CIFAR-100F | ImageNet1k |
| 1 5 | 15.31M | ||||||
| 2 5 | 16.94M | ||||||
| 7 10 | 16.94M | ||||||
| 1 3 | 18.56M | ||||||
| 2 4 | 18.56M | ||||||
| 3 5 | 18.56M | ||||||
| 1 2, 3 4 | 18.56M | ||||||
| 1 2, 4 5 | 18.56M | ||||||
| 0 1 | 20.43M | ||||||
| 1 2 | 20.43M | ||||||
| 2 3 | 20.43M | ||||||
| 3 4 | 20.43M | ||||||
| 4 5 | 20.43M | ||||||
| 5 6 | 20.43M | ||||||
| 6 7 | 20.43M | ||||||
| 7 8 | 20.43M | ||||||
| 8 9 | 20.43M | ||||||
| 9 10 | 20.43M | ||||||
| 10 11 | 20.43M | ||||||
| original | 22.06M |
| Approx. | Params. | MNIST | F-MNIST | CIFAR-10 | CIFAR-100C | CIFAR-100F | ImageNet1k |
| 1 5 | 15.31M | ||||||
| 2 5 | 16.94M | ||||||
| 7 10 | 16.94M | ||||||
| 1 3 | 18.56M | ||||||
| 2 4 | 18.56M | ||||||
| 3 5 | 18.56M | ||||||
| 1 2, 3 4 | 18.56M | ||||||
| 1 2, 4 5 | 18.56M | ||||||
| 0 1 | 20.43M | ||||||
| 1 2 | 20.43M | ||||||
| 2 3 | 20.43M | ||||||
| 3 4 | 20.43M | ||||||
| 4 5 | 20.43M | ||||||
| 5 6 | 20.43M | ||||||
| 6 7 | 20.43M | ||||||
| 7 8 | 20.43M | ||||||
| 8 9 | 20.43M | ||||||
| 9 10 | 20.43M | ||||||
| 10 11 | 20.43M | ||||||
| original | 22.06M |
| Approx. | Params. | MNIST | F-MNIST | CIFAR-10 | CIFAR-100C | CIFAR-100F | ImageNet1k |
| 1 5 | 15.31M | ||||||
| 2 5 | 16.94M | ||||||
| 7 10 | 16.94M | ||||||
| 1 3 | 18.56M | ||||||
| 2 4 | 18.56M | ||||||
| 3 5 | 18.56M | ||||||
| 1 2, 3 4 | 18.56M | ||||||
| 1 2, 4 5 | 18.56M | ||||||
| 0 1 | 20.43M | ||||||
| 1 2 | 20.43M | ||||||
| 2 3 | 20.43M | ||||||
| 3 4 | 20.43M | ||||||
| 4 5 | 20.43M | ||||||
| 5 6 | 20.43M | ||||||
| 6 7 | 20.43M | ||||||
| 7 8 | 20.43M | ||||||
| 8 9 | 20.43M | ||||||
| 9 10 | 20.43M | ||||||
| 10 11 | 20.43M | ||||||
| original | 22.06M |
| Accuracy | ||||||
| Approx. | Params. | MNIST | F-MNIST | CIFAR-10 | CIFAR-100C | CIFAR-100F |
| 1 5 | -25.99M | |||||
| 2 5 | -19.49M | |||||
| 1 3 | -13M | |||||
| 3 5 | -13M | |||||
| 2 4 | -13M | |||||
| 8 10 | -13M | |||||
| 9 11 | -13M | |||||
| 3 4 | -6.5M | |||||
| 4 5 | 6.5M | |||||
| 0 1 | -6.5M | |||||
| 1 2 | -6.5M | |||||
| 2 3 | -6.5M | |||||
| - | 86.39M | |||||
| Accuracy | |||||||
| Approx. | Params. | MNIST | F-MNIST | CIFAR-10 | CIFAR-100C | CIFAR-100F | ImageNet1k |
| 112 | -127.02M | ||||||
| 04 | -46.19M | ||||||
| 15 | -46.19M | ||||||
| 25 | -34.64M | ||||||
| 710 | -34.64M | ||||||
| 02, 34 | -34.64M | ||||||
| 12, 34 | -23.09M | ||||||
| 13 | -23.09M | ||||||
| 24 | -23.09M | ||||||
| 35 | -23.09M | ||||||
| 57 | -23.09M | ||||||
| 01 | -11.55M | ||||||
| 12 | -11.55M | ||||||
| 23 | -11.55M | ||||||
| 34 | -11.55M | ||||||
| 45 | -11.55M | ||||||
| 56 | -11.55M | ||||||
| 67 | -11.55M | ||||||
| 78 | -11.55M | ||||||
| 89 | -11.55M | ||||||
| 910 | -11.55M | ||||||
| 1011 | -11.55M | ||||||
| 1112 | -11.55M | ||||||
| 1213 | -11.55M | ||||||
| 1314 | -11.55M | ||||||
| 1415 | -11.55M | ||||||
| 1516 | -11.55M | ||||||
| 1617 | -11.55M | ||||||
| 1718 | -11.55M | ||||||
| 1819 | -11.55M | ||||||
| 1920 | -11.55M | ||||||
| 2021 | -11.55M | ||||||
| 2122 | -11.55M | ||||||
| 2223 | -11.55M | ||||||
| original | 304.35M | ||||||
A.2.4 Transformation Generalization
In this section, we extend the analysis to evaluate the model’s ability to generalize transformations across datasets to different approximation blocks and models (i.e., DiNO-S).
| Accuracy | ||||||
| Encoder | Approx. | Source | MNIST | CIFAR-10 | CIFAR-100C | CIFAR-100F |
| ViT-S | 2 3 | MNIST | ||||
| CIFAR-10 | ||||||
| CIFAR-100 | ||||||
| 3 4 | MNIST | |||||
| CIFAR-10 | ||||||
| CIFAR-100 | ||||||
| 4 5 | MNIST | |||||
| CIFAR-10 | ||||||
| CIFAR-100 | ||||||
| 9 10 | MNIST | |||||
| CIFAR-10 | ||||||
| CIFAR-100 | ||||||
| 1 3 | MNIST | |||||
| CIFAR-10 | ||||||
| CIFAR-100 | ||||||
| 3 5 | MNIST | |||||
| CIFAR-10 | ||||||
| CIFAR-100 | ||||||
| 2 4 | MNIST | |||||
| CIFAR-10 | ||||||
| CIFAR-100 | ||||||
| 8 10 | MNIST | |||||
| CIFAR-10 | ||||||
| CIFAR-100 | ||||||
| 9 11 | MNIST | |||||
| CIFAR-10 | ||||||
| CIFAR-100 | ||||||
| 2 5 | MNIST | |||||
| CIFAR-10 | ||||||
| CIFAR-100 | ||||||
| 7 10 | MNIST | |||||
| CIFAR-10 | ||||||
| CIFAR-100 | ||||||
| 1 5 | MNIST | |||||
| CIFAR-10 | ||||||
| CIFAR-100 | ||||||
| Accuracy | ||||||
| Encoder | Approx. | Source | MNIST | CIFAR-10 | CIFAR-100C | CIFAR-100F |
| DiNO-S | 2 3 | MNIST | ||||
| CIFAR-10 | ||||||
| CIFAR-100 | ||||||
| 3 4 | MNIST | |||||
| CIFAR-10 | ||||||
| CIFAR-100 | ||||||
| 4 5 | MNIST | |||||
| CIFAR-10 | ||||||
| CIFAR-100 | ||||||
| 9 10 | MNIST | |||||
| CIFAR-10 | ||||||
| CIFAR-100 | ||||||
| 1 3 | MNIST | |||||
| CIFAR-10 | ||||||
| CIFAR-100 | ||||||
| 3 5 | MNIST | |||||
| CIFAR-10 | ||||||
| CIFAR-100 | ||||||
| 2 4 | MNIST | |||||
| CIFAR-10 | ||||||
| CIFAR-100 | ||||||
| 8 10 | MNIST | |||||
| CIFAR-10 | ||||||
| CIFAR-100 | ||||||
| 9 11 | MNIST | |||||
| CIFAR-10 | ||||||
| CIFAR-100 | ||||||
| 2 5 | MNIST | |||||
| CIFAR-10 | ||||||
| CIFAR-100 | ||||||
| 7 10 | MNIST | |||||
| CIFAR-10 | ||||||
| CIFAR-100 | ||||||
| 1 5 | MNIST | |||||
| CIFAR-10 | ||||||
| CIFAR-100 | ||||||
A.2.5 Comparison
In this section, we provide additional results for the SkipAt baseline.
| Accuracy | |||||
| Skip | MNIST | F-MNIST | CIFAR-10 | CIFAR-100C | CIFAR-100F |
| 1 5 | |||||
| 2 5 | |||||
| 7 10 | |||||
| 1 3 | |||||
| 3 5 | |||||
| 2 4 | |||||
| 8 10 | |||||
| 9 11 | |||||
| 0 1 | |||||
| 1 2 | |||||
| 2 3 | |||||
| 3 4 | |||||
| 4 5 | |||||
| 5 6 | |||||
| 6 7 | |||||
| 7 8 | |||||
| 8 9 | |||||
| 9 10 | |||||
| 10 11 | |||||
| original | |||||
A.2.6 Analysis of misclassifications
In this section, we examine changes in per-class accuracy and misclassification patterns. As shown in Figure 12, models behave differently at block approximations. DiNO-S remains remarkably stable across blocks and classes, with the only degradation appearing for classes dog (when approximating blocks 10 or 11) and deer (for block 10 approximation). ViT-S shows a similar drop for class dog on its final block. Instead, the most noticeable hit occurs for class cat when the earlier blocks are approximated. For DEiT-S, several mid-to-late block approximations improve accuracy for various classes, whereas the very first block causes a clear relative decline in nearly every class. These observations suggest strategies like preferring late-block approximation for DEiT-S, or reserving extra samples for the linear transformation in order to recover the accuracy of difficult classes for the model.
In order to further investigate how the predictions change while approximating blocks, we plot the difference in the normalized confusion matrix before and after the approximation. In Figure 13, we show the delta confusion matrix for DEiT-S on CIFAR-100C. Also, here we can see how approximating the very first block makes the model puzzling and lose per-class accuracy (i.e., negative delta along the diagonal). On the other hand, approximating the last block acts as a regularizer, resulting in an overall gain in the per-class accuracy and, as a consequence, fewer misclassifications (negative deltas off-diagonal). This supports results shown in Figures 12 and 1.
Additionally, Figure 14 shows representative CIFAR-10 images that become misclassified after approximating a block of ViT-S. The patterns we observe mirror the trends in Figures 12 and 13: when approximating earlier blocks, we observe many images belonging to class cat to be misclassified. Instead, when approximating later blocks, we observe images of the class dog to be misclassified. Together, these qualitative examples show that understanding these block-specific vulnerabilities allows us to steer the approximation procedure, informing choices about which blocks to approximate based on the observed impact on the final model’s class-wise performance.

### v6

@@
-Foundation Models on a Budget: Approximating Blocks in Large Vision Models
+: Transformer Optimization using Adaptive and Simple Transformations
 Abstract
-Foundation Models have shown impressive performance in various tasks and domains, yet they require massive computational resources, raising concerns about accessibility and sustainability.
-Previous attempts to reduce foundation model size fall short of fully addressing the problem, as they end up increasing computational load through additional training steps.
-Recent works reveal that deep neural networks exhibit internal representation similarities.
-While inter-network similarities have enabled techniques such as model stitching and merging, intra-network similarities remain underexplored for improving efficiency.
-In this paper, we propose Transformer Blocks Approximation (TBA), a novel method that leverages intra-network similarities to identify and approximate transformer blocks in large vision models.
-TBA replaces these blocks using lightweight, closed-form transformations, without retraining or fine-tuning the rest of the model.
-The proposed method reduces the number of parameters while having minimal impact on the downstream task.
-We validate the effectiveness and generalizability of TBA through extensive experiments across multiple datasets (e.g., Imagenet-1k and CIFAR100) and state-of-the-art pretrained vision models (e.g, ViT, DiNO-v2, and DEiT).
+Foundation models achieve State-of-the-art (SOTA) performance across different tasks, but their size and computational demands raise concerns about accessibility and sustainability.
+Existing efficiency methods often require additional retraining or fine-tuning, limiting their practicality.
+Recent findings suggest that deep neural networks exhibit internal representation similarities.
+While such similarities across different models have been exploited for enabling techniques such as model stitching and merging, intra-network redundancy remains underexplored as a source for efficiency gains.
+In this paper, we introduce Transformer Optimization using Adaptive and Simple Transformations (TOAST), a framework that exploits these redundancies to approximate entire transformer blocks with lightweight closed-form mappings, such as linear transformation or even the identity, without any additional training.
+Across SOTA pretrained vision models (e.g., ViT, DINOv2, DeiT) and datasets ranging from MNIST to ImageNet-1k, TOAST reduces parameters and computation while preserving, and in some cases improving, downstream performance.
+These results show that large portions of transformer depth can be replaced by trivial functions, opening a new perspective on efficient foundation models.
 1 Introduction
-As Neural Networks (NNs) grow in size and complexity, their demand for computational resources has become a significant bottleneck.
-Despite the impressive performance of large models, they often come with substantial trade-offs, such as increased memory and power consumption.
-This has led to a growing interest in methods that can reduce model complexity without sacrificing performance.
-However, most existing approaches trying to mitigate these challenges either demand additional training steps or result in substantial performance degradation.
-Recent research showed that there exist similarities between representation within and between NNs.
-The observation that many layers or components within these networks perform similar functions or produce highly correlated outputs suggests that some blocks may be approximated, highlighting opportunities for network simplification.
-In this paper, we investigate two key research questions: (i) how to effectively identify blocks within neural networks that yield similar representations, and (ii) how to approximate these blocks with minimal impact on overall performance without any additional training steps.
-To address the first question, we provide an extensive analysis showing that the Mean Squared Error (MSE) effectively identifies such blocks, suggesting that their approximation should have a negligible impact.
-For the second question, we propose Transformer Blocks Approximation (TBA), a novel method that leverages similarities between representations to approximate transformer blocks using lightweight transformations, such as linear mappings.
-Once the blocks that have minimal impact on model functionality are identified, TBA replaces them with a simpler transformation.
-This approximation strategy allows for reducing the overall parameter count, ensuring minimal impact on downstream task performance.
+As Neural Networks (NNs) continue to grow in size and complexity, their demand for computational resources has become a critical bottleneck.
+While larger models consistently achieve SOTA performance, this comes at the cost of substantial memory usage and power consumption, limiting their accessibility and deployment.
+This challenge is for instance most relevant in on-device scenarios, where saving memory, latency, and energy, even by little margins, is critical (Pan et al., 2022; Li et al., 2022).
+This has motivated a growing body of work on reducing model complexity.
+However, most existing approaches either require additional, resource-intensive training phases or lead to significant drops in accuracy.
+Recent studies reveal that there exists strong representational similarities both within and between NNs.
+In other words, when focusing on intra-network similarities, different blocks often perform overlapping functions or produce highly correlated outputs.
+This redundancy suggests an opportunity: instead of retraining or pruning, can we approximate these blocks with simpler transformations?
+To address this question, we propose Transformer Optimization using Adaptive and Simple Transformations (TOAST), a novel framework that exploits block-level representational redundancy to replace transformer blocks with lightweight transformations.
+By doing so, TOAST reduces parameter count and computational cost, while maintaining (and in some cases even improving) downstream task performance.
+Crucially, our method is training-free, making it simple, efficient, and widely applicable, even in resource-constrained scenarios such as deployment on edge devices, where even the smallest available models may exceed memory or power budgets.
 Our main contributions are as follows:
-We introduce TBA, a general framework for identifying and approximating transformer blocks that produce similar outputs, using simpler transformations (e.g., linear), reducing model parameters and complexity with minimal to no impact on the downstream task performance (Figure 1).
+We propose TOAST, a simple yet effective framework that replaces transformer blocks with lightweight transformations (e.g., linear maps or even the identity), significantly reducing parameters and computational cost while preserving downstream performance (Figure˜1).
 2 Related work
-Measuring similarities A range of metrics have been introduced to assess the similarity between latent spaces generated by different NNs (Klabunde et al., 2023; Ballester et al., 2023).
+Measuring Similarities
+A range of metrics have been introduced to assess the similarity between latent spaces generated by different NNs (Klabunde et al., 2023; Ballester et al., 2023).
 One established approach is Canonical Correlation Analysis (CCA) (Hotelling, 1992), known for its invariance to linear transformations.
@@
 (2024) leverages the principles of spectral geometry to model and analyze the relationships between distinct latent spaces.
-Leveraging similarities Valeriani et al.
+Leveraging Similarities
+Valeriani et al.
 (2024) examine the intrinsic dimensionality and neighbor compositions of representations in transformer models.
@@
 (2024) demonstrate that representations learned by distinct NNs can be aligned using simple transformations.
-Architectural efficiency While large-scale models with billions or even trillions of parameters continue to achieve state-of-the-art performance, their growth comes with trade-offs, such as slower inference times and significantly higher computational costs.
+Architectural Efficiency
+While large-scale models with billions or even trillions of parameters continue to achieve state-of-the-art performance, their growth comes with trade-offs, such as slower inference times and significantly higher computational costs.
 Improving the efficiency of Deep Neural Network (DNN) has been a long-standing area of research.
@@
 (2023) shows that for CNNs, this additional training step can sometimes be avoided.
-Unlike other methods, TBA leverages intra-network similarities to reduce foundation model complexity without the need for additional training steps while maintaining competitive performance.
-3 Transformer Blocks Approximation
-The core idea of the proposed approach is to detect similar representations within NNs and replace the corresponding blocks with simple transformations.
-In the context of this work, a "block" refers to a sequence of layers including multi-head self-attention, normalization, and feed-forward layers, functioning as a cohesive unit.
-Instead of executing the entire NN, using transformer block approximation (TBA), we approximate these blocks, reducing computational complexity while preserving the network’s functionality.
-To provide a clear understanding, Figure 1 presents an overview of our method.
-Identifying similar representations We hypothesize that certain foundation models may contain blocks that produce similar representations.
-To quantify these similarities, we employ the MSE.
-A low MSE between the output of an earlier block and a later block indicates that their respective representations are highly similar.
-This suggests that the intervening sequence of blocks contributes minimally to transforming the representation, highlighting potential redundancy.
-Let represent the total number of blocks in the model.
-Let denote the output representation (a vector of dimension ) of block for a single input, where .
-For any two blocks, a starting block and an ending block (where ), we compute their output representations and .
-For each , where is a subset of points sampled uniformly at random from , defines the input to the corresponding layer .
-The MSE between these representations is defined as:
-By systematically evaluating the MSE for various pairs of blocks , we can identify those being good candidates for approximation.
-This allows for targeted complexity reduction without substantially altering the network’s internal representations or downstream-task performance at a very low computational cost.
-Approximating transformer blocks Once two blocks and are identified as having highly similar output representations, our goal is to replace the intermediate blocks with a single, lightweight transformation that maps the output of block directly to an approximation of the output of block .
+Unlike other methods, TOAST leverages intra-network similarities to reduce vision transformers complexity without the need for additional training steps while maintaining competitive performance.
+3 Blocks Approximation
+The central idea of our approach is that it is possible to leverage representation similarities within transformer-based architectures to replace entire attention blocks with simpler transformations.
+In this work, a block refers to a sequence of layers including multi-head self-attention, normalization, and feed-forward layers, that function together as a cohesive unit.
+By replacing these blocks with simpler transformations, we can reduce the computational complexity of the network while maintaining its core functionality.
+Approximating Transformer Blocks
+Given two blocks and , our goal is to replace the intermediate blocks with a single, lightweight transformation that maps the output of block directly to an approximation of the output of block .
 This approach allows us to skip the computation of blocks , effectively reducing the overall computational costs.
 This approximation can be repeated for multiple, non-overlapping blocks, i.e., blocks and with .
-Let and represent the output representations from block and respectively, for the data points in . Our objective is to find a transformation such that:
-In this work, we consider to be a linear transformation that can be estimated by minimizing the squared error between the transformed output and the actual , which can be solved using least squares:
+An overview of the method is provided in Figure˜1.
+Let and represent the output representations from block and respectively, for the data points in , sampled uniformly at random from the full training dataset .
+Our objective is to find a transformation such that:
+In this work, we consider to be the identity or a linear transformation . We can compute the linear transformation by minimizing the squared error between the transformed output and the actual :
 This optimization problem allows for a closed-form solution that efficiently computes the optimal transformation .
 The solution bypasses the computation of all layers between any two blocks and , replacing them with .
-This approximation significantly reduces computational complexity by replacing one or more transformer blocks, comprising multi-head self-attention and feed-forward layers, with a low-cost linear transformation.
-To sum up, our approach consists of two main stages:
-1.
-Identification: We use MSE to identify blocks that produce similar representations.
-2.
-Approximation: We compute an efficient linear approximation using a transformation matrix to bypass the previously identified blocks.
-This procedure reduces parameter count with minimal impact on internal representations, as shown in Figures 9, 7, 10, 11 and 8, and maintains compatibility with any downstream classifiers, achieving significant compression as shown Tables 1, 8, 9, 10, 11, 12 and 13.
+This approximation reduces computational complexity while minimally affecting internal representations, as illustrated in Figures˜9, 7, 10, 11 and 8, and preserves compatibility with downstream classifiers, achieving significant compression as shown in Tables˜1, 2, 3, 7, 8, 9 and 10.
+Patterns of Similarity between Transformer Blocks
+Inspired by existing results Kornblith et al.
+(2019); Nguyen et al.
+(2020), and in particular Venkataramanan et al.
+(2024), which show that multi-head attention modules exhibit similarity in learned representations, we investigate whether pre-trained foundation models contain entire blocks that produce highly similar representations.
+To investigate this hypothesis, we adopt the CKA, where a high score between the outputs of two blocks indicates that their respective representations are highly similar.
+This, in turn, suggests that the intermediate blocks contribute minimally to transforming the representation, highlighting potential redundancy in the model.
+After centering the representations, i.e., and , by subtracting their respective means, the CKA is computed as:
+Here, denotes the Frobenius norm. A CKA value closer to 1 indicates a higher degree of similarity between the representations of the two blocks.
+By systematically evaluating the CKA across pairs of blocks , it is possible to distinguish patterns of similarities between different blocks.
+These patterns indicate opportunities for targeted complexity reduction, enabling the approximation of redundant blocks without substantially altering internal representations or downstream task performance.
 4 Experiments
-This section begins with an analysis of block-level similarities within pre-trained vision foundation models (Section 4.1), and we investigate how TBA impacts the representations of these latent spaces.
-Following this, Section 4.2 provides a comprehensive quantitative evaluation of TBA’s performance on various image classification tasks, including standard, zero-shot, and cross-dataset generalization scenarios.
-Lastly, in Section 4.3 we compare TBA against existing baselines, including an ablation study on transformation complexity, further validating the effectiveness of the proposed approach.
+In this section, we first analyze the similarities between different transformer blocks to motivate their approximation using simple transformations.
+We then present comprehensive results on image classification across various models and datasets to demonstrate the effectiveness and efficiency of the proposed method.
+Beyond these core results, we further study the robustness of TOAST through ablations on the number of samples required for approximation and the choice of translator architecture.
+Overall, our findings show that TOAST achieves strong performance while producing lighter and faster models.
+Due to space constraints, additional results on zero-shot image classification, as well as further qualitative and quantitative analyses, are provided in the Appendix (Sections˜A.2.3 and A.2.2).
 4.1 Latent analysis
-Block similarities We investigate similarities emerging in latent representations using three distinct transformer-based models with different dimensionality: ViT-S, DiNO-B, and DEiT-S, and five datasets: CIFAR-10, CIFAR-100, MNIST, F-MNIST, and ImageNet1k.
-Details on the models and datasets are provided in Table 6 and Table 7.
-Figure 2 presents the MSE matrices between blocks of different models when using different datasets.
-These similarities are calculated using only a small subset of the training data (i.e., 500 samples).
-Results reveal that while the patterns of similarity vary across models, they remain consistent across different datasets.
-This suggests that the similarity structure between computational blocks is predominantly influenced by the model rather than the dataset used.
-This finding aligns with observations from Nguyen et al.
-(2020), where DNNs trained from scratch tend to exhibit a distinctive “block structure” in their representations, linked to model overparameterization.
-Our results extend this observation to pretrained foundation models, where we show that such a structure is primarily influenced by the model.
-We conduct the analysis using the mean over the tokens, but in Section A.2.1 we show that this also applies when using only the CLS token from each block.
-Finally, in Figure 6, we present an ablation study to investigate the influence of the similarity metric.
-We correlate various metrics with downstream task performance using ViT-S and ImageNet1k.
-Our findings reveal that using simple metrics such as MSE or cosine similarity to identify block-level similarities exhibits a strong correlation with final accuracy.
-These results underscore the flexibility of our proposed method, indicating its compatibility with a range of similarity or distance functions that can capture inter-block relationships.
-Effects of TBA on latent representations After identifying blocks producing similar representations, we analyze the impact on the latent representation output by the last block when approximating using TBA.
-These blocks are approximated using a shared linear transformation applied across all tokens based on a subset of 3,000 training samples.
-For consistency, we employed the same models and datasets used for the results in Figure 2.
-We compute the MSE between the representations in the last block of the original and the TBA model when approximating the block.
-As illustrated in Figure 3, in most cases, the MSE increases as the block depth increases.
-Original TBA
+In this section we investigate similarities in the latent representations of DiNO-B and DEiT-S on five datasets: CIFAR-10, CIFAR-100, MNIST, F-MNIST, and ImageNet1k.
+We compute the CKA similarity metric averaged over all tokens from a small subset of 100 training samples, which is sufficient to reveal block-level similarity patterns while remaining computationally efficient.
+Additional results with other pretrained vision transformers (ViT-T, ViT-S, DiNO-S, ViT-B) are provided in Section˜A.2.1, showing consistent patterns for each model across different datasets.
+Do vision transformer models exhibit block-wise similarity patterns?
+The results in Figure˜2 reveal that while the similarity patterns differ across models, they remain largely consistent for the same model across different datasets.
+This suggests that the similarity structure between computational blocks is predominantly influenced by the model itself.
+Although the general similarity pattern remains the same, the differences in CKA values become more pronounced (i.e., the block structure becomes more evident) as the complexity of the dataset increases (e.g., from MNIST to ImageNet1k).
+These finding aligns with observations from Nguyen et al.
+(2020), where DNN trained from scratch exhibit a distinctive "block structure" in their representations, which is linked to model overparameterization.
+Our results extend this observation to vision pre-trained foundation models, showing that such a structure is primarily an intrinsic property of the model.
+Moreover, these consistent block-wise patterns indicate potential targets for approximation, suggesting that entire blocks may be replaced with simpler transformations without substantially altering the model’s internal representations.
+How does TOAST affect latent representations?
+We next analyze the impact of the proposed transformations on the final block’s latent representations, which are used for downstream classification.
+We approximate these blocks using a shared linear transformation applied across all tokens, estimated on a subset of 500 training samples.
+For consistency, we use the same models and datasets as in Figure˜2.
+Original
+TOAST
 DiNO-B
 DEiT-S
-This suggests that approximating the final blocks would lead to significant changes in the final representations, indicating their critical role.
-However, in the case of DEiT-S, the trend is different.
-Here, the MSE decreases toward the final blocks.
-This is also shown in the similarity matrices illustrated in Figure 2, where the final blocks have a high similarity.
-These findings reinforce the intuition behind the use of MSE, demonstrating a correlation between the metric and the final representation similarity when approximating blocks.
-In some instances, such as with the MNIST dataset, the MSE scores remain relatively consistent across blocks, indicating that the representations are largely similar to one another.
-However, for more complex datasets like ImageNet1k, the representations in the final or the first blocks become increasingly dissimilar, making it advantageous to approximate intermediate blocks.
-This suggests that the similarities are primarily driven by the model but also guided by the complexity of the dataset, allowing for targeted approximations that reduce model parameters and complexity without significantly compromising performance.
-To further investigate the effect of TBA on the latent representations, we visualize, using Principal Component Analysis (PCA), the representations of the TBA model when approximating the final block and its original representation (see Figure 4).
-The plot shows the representations generated using the DiNO-B and DEiT-S pretrained encoders on F-MNIST.
-Colors represent classes.
-The visualization highlights that approximating the final block of DiNO-B results in noticeable deviations from the original representations, while in DEiT-S the approximated representations remain similar to the original ones.
-This observation aligns with the results from Figure 3, where approximating blocks can lead to significant changes in representations.
-For complete results, refer to Section A.2.2.
-4.2 Transformer block approximation performance
-We perform image classification using various pretrained models with different dimensionalities (ViT-T, ViT-S, ViT-L, ViT-B, DiNO-S, DiNO-B, and DEiT-S) and datasets (MNIST, F-MNIST, CIFAR-10, CIFAR-100C, CIFAR-100F, and ImageNet1k), keeping all models frozen.
-Block approximations are computed using a shared linear transformation applied across all tokens, based on
-| Accuracy | |||||
-| Approx. | Params. | CIFAR-10 | CIFAR-100F | ImageNet1k | |
-| ViT-S | 1 5 | -6.51M | (-11.09) | (-27.35) | (-39.86) |
-| 2 5 | -4.88M | (-4.88) | (-14.42) | (-17.56) | |
-| 7 10 | -4.88M | (-8.69) | (-23.89) | (-51.14) | |
-| 2 4 | -3.26M | (-2.95) | (-8.67) | (-6.80) | |
-| 9 11 | -3.26M | (-5.00) | (-14.71) | (-35.54) | |
-| 1 2, 4 5 | -3.26M | (-3.54) | (-11.10) | (-12.06) | |
-| 0 1 | -1.63M | (-2.09) | (-6.15) | (-11.10) | |
-| 3 4 | -1.63M | (-0.66) | (-2.47) | (-2.51) | |
-| 9 10 | -1.63M | (-1.15) | (-6.19) | (-15.58) | |
-| 10 11 | -1.63M | (-1.67) | (-4.97) | (-12.73) | |
-| original | 21.82M | ||||
-| DiNO-B | 1 5 | -26.5M | (-11.28) | (-28.62) | (-58.57) |
-| 2 5 | -19.5M | (-5.87) | (-19.50) | (-28.26) | |
-| 7 10 | -19.5M | (-4.46) | (-14.02) | (-35.88) | |
-| 2 4 | -13M | (-1.49) | (-6.62) | (-7.70) | |
-| 9 11 | -13M | (-10.76) | (-14.36) | (-27.74) | |
-| 1 2, 4 5 | -13M | (-1.47) | (-6.11) | (-7.79) | |
-| 0 1 | -6.5M | (-0.05) | (-1.01) | (-1.00) | |
-| 3 4 | -6.5M | (-0.38) | (-2.78) | (-1.62) | |
-| 9 10 | -6.5M | (-0.78) | (-3.26) | (-10.09) | |
-| 10 11 | -6.5M | (-1.01) | (-4.23) | (-13.83) | |
-| original | 86.58M | ||||
-| DEiT-S | 1 5 | -6.51M | (-13.55) | (-28.90) | (-40.88) |
-| 2 5 | -4.88M | (-5.41) | (-14.87) | (-16.05) | |
-| 7 10 | -4.88M | (-1.61) | (-2.78) | (-22.24) | |
-| 2 4 | -3.26M | (-1.85) | (-6.37) | (-5.32) | |
-| 9 11 | -3.26M | (+0.30) | (+1.11) | (-5.39) | |
-| 1 2, 4 5 | -3.26M | (-5.74) | (-13.31) | (-10.68) | |
-| 0 1 | -1.63M | (-6.21) | (-12.91) | (-15.38) | |
-| 3 4 | -1.63M | (-0.14) | (-1.24) | (-1.22) | |
-| 9 10 | -1.63M | (+0.30) | (+0.86) | (-2.15) | |
-| 10 11 | -1.63M | (+0.49) | (+1.15) | (+0.05) | |
-| original | 21.82M | ||||
-a subset of 3,000 training samples.
-Then, we train a single linear layer using the Adam optimizer (learning rate ) for 5 epochs, with 3 seeds and a batch size of 256.
-Details regarding the models, datasets, computational resources, and software tools are provided in Table 6, Table 7, Section A.1.4, and Section A.1.3, respectively.
-Image classification As presented in Table 1, the proposed TBA method successfully reduces model size while maintaining, and in some cases even improving, image classification performance.
-These downstream task results support the findings from the analysis in Section 4.1.
-Specifically, as discussed in Section 4.1 and illustrated in Figures 3 and 4, approximating the final block of DEiT-S yields a final representation highly similar to the original.
-This similarity suggests such blocks are optimal candidates for approximation, a conclusion further validated by Table 1.
-The table shows that even when approximating multiple consecutive blocks (e.g., 911), we can reduce parameters while achieving performance superior to the original model.
-Overall, the consistent or enhanced performance indicates that a linear transformation is sufficient to approximate these transformer blocks, thereby effectively reducing model parameters.
-It’s important to note that this transformation is shared across all tokens, with no additional training steps required.
-Additional results are presented in Tables 8, 9, 10, 11, 12 and 13.
-To assess the scalability of our method, we extended our evaluation to ViT-L.
-While our approach reduces its parameters, ViT-L still imposes considerable computational demands, necessitating high-performance GPU infrastructure.
-This finding underscores that while effective for substantial compression, our method’s impact on making extremely large models practical for resource-constrained environments is limited by their initial scale.
-Finally, in Section A.2.6, we further analyze model behavior after block approximations on the final image-classification task, complementing the quantitative results with qualitative insights into the approximation procedure.
-Zero-shot image classification To further assess the effectiveness of our approach, we evaluate
-| Params. | Approx. | Accuracy | |
-| -6.49M | 0 1 | 57.93 | - |
-| 1 2 | 64.20 | - | |
-| 2 3 | 66.35 | -5.51% | |
-| 3 4 | 64.65 | - | |
-| 4 5 | 64.86 | - | |
-| 5 6 | 58.05 | - | |
-| 6 7 | 61.56 | - | |
-| 7 8 | 58.53 | - | |
-| 8 9 | 52.32 | - | |
-| 9 10 | 59.21 | - | |
-| 10 11 | 22.64 | - | |
-| 149.07M | original | 70.21 | – |
-TBA in a zero-shot image classification setting.
-This evaluation utilizes the OpenCLIP-ViT-B model (Radford et al., 2021), which was pretrained on LAION-2B Schuhmann et al.
-(2022), with ImageNet1k serving as the downstream evaluation dataset.
-As in previous experiments, the model remains frozen, and block approximations are computed using a shared linear transformation applied across all tokens, based on a subset of 3,000 training samples.
-Importantly, we apply these approximations only to the vision encoder, leaving the text encoder unchanged.
-We follow the standard ImageNet1k prompt templates.
-As shown in Table 2, when approximating earlier vision blocks (e.g., 1 2 or 2 3), the model still achieves competitive results, but slightly reduces the number of parameters.
-The analysis is conducted only on the base version, as larger versions (e.g., OpenCLIP-ViT-L or OpenCLIP-ViT-H) contain too many parameters and are thus beyond the scope of this paper.
-Details on the models are provided in Table 6.
-Transformation generalization We further evaluate the model’s ability to generalize transformations across datasets.
-Specifically, we analyze the capability of applying a linear transformation learned on one dataset (source) to another (target), using the same underlying pretrained architecture and without any subsequent training or fine-tuning of the approximated model.
-Table 3 reports accuracy scores for ViT-S and DiNO-S on MNIST, CIFAR-10, and CIFAR-100, where the "Source" column indicates the source dataset used to compute the transformation.
-The results demonstrate that a single, token-shared linear transformation can generalize effectively across different target datasets, with the exception of MNIST, which may be too simplistic to yield broadly transferable representations.
-This cross-dataset generalization is particularly valuable in scenarios where fine-tuning is infeasible, such as in privacy-sensitive domains like healthcare or finance.
-In these contexts, regulatory constraints often restrict training on local data, yet leveraging existing pretrained models remains crucial.
-Our findings suggest that lightweight adaptation via shared linear transformations offers a promising path for improving model performance on custom datasets without requiring additional training.
-Additional results are provided in Tables 14 and 15.
-| Accuracy | ||||
-| Approx. | Source | MNIST | CIFAR-10 | CIFAR-100C |
-| 3 4 | MNIST | |||
-| CIFAR-10 | ||||
-| CIFAR-100F | ||||
-| 3 5 | MNIST | |||
-| CIFAR-10 | ||||
-| CIFAR-100F | ||||
-4.3 Comparison
-Direct comparisons for TBA are limited, as existing approaches either fine-tune pretrained models or train from scratch.
-In contrast, TBA maintains the foundation model frozen.
-We include two types of comparisons: a variant of SkipAt, a method proposed in Venkataramanan et al.
-(2024), and an ablation study on transformation complexity.
-SkipAt identity Venkataramanan et al.
-(2024) explores the effect of skipping self-attention or multi-head self-attention by inserting identity functions in pretrained models, without retraining.
-We extend this idea by skipping entire blocks, not just attention components.
-As shown in Table 4, TBA consistently outperforms this SkipAt Identity baseline on CIFAR-10 and CIFAR-100F.
-The accuracy improvements confirm that approximating is more effective than skipping them.
-Results with other datasets and models can be found in Section A.2.5.
-| CIFAR-10 | CIFAR-100F | |||
-| Approx. | SkipAt | TBA | SkipAt | TBA |
-| 1 5 | ||||
-| 2 5 | ||||
-| 7 10 | ||||
-| 1 3 | ||||
-| 2 4 | ||||
-| 9 11 | ||||
-| 2 3 | ||||
-| 9 10 | ||||
-| original | ||||
-Trained approximators We also evaluate whether more complex approximations improve performance.
-Specifically, we compare TBA to deeper MultiLayer Perceptron (MLP)-based translators (MLP and Res-MLP), which are trained for 300 steps using Adam with a learning rate of 1e-3.
-As shown in Table 5, employing a simple linear transformation to approximate blocks is the optimal choice in most cases.
-While deeper translators occasionally yield slightly higher accuracy (e.g., 15), they incur significant training overhead and require gradient-based optimization.
-In contrast, TBA is entirely training-free and operates in closed form; no backpropagation or updates are required.
-This highlights the efficiency–performance trade-off and supports the design choice behind TBA.
-Refer to Section A.1.2 for approximator details.
-| Approx. | TBA | MLP | Res-MLP |
-| 1 5 | |||
-| 2 5 | |||
-| 7 10 | |||
-| 1 3 | |||
-| 2 4 | |||
-| 9 11 | |||
-| 3 4 | |||
-| 9 10 |
+To quantify the effect of the approximation, we compute the CKA similarity between the final block representations of the original and the TOAST-approximated model for each block using its preceding block as input.
+As shown in Figure˜4, the model-specific similarity patterns re-emerge after approximation.
+The plots highlight more specific trends.
+Approximating blocks is easier on simpler tasks (e.g., image classification on MNIST or F-MNIST), yielding representations that closely match the originals, whereas on more complex datasets (e.g., ImageNet1k or CIFAR-100), the approximated representations deviate more from the original ones.
+Furthermore, the final blocks of DEiT-S exhibit high similarity, suggesting that approximating these layers preserves the final representations, while earlier blocks remain more critical.
+To provide a more intuitive view, Figure˜4 visualizes the final-layer representations using Principal Component Analysis (PCA).
+We compare the original representations with those obtained after approximating the final block (10 11) using TOAST on F-MNIST, with colors indicating the 10 classes.
+The visualization confirms that approximating the final block of DiNO-B results in noticeable deviations from the original representations, whereas for DEiT-S the approximated representations remain highly similar.
+These observations align with the CKA analysis in Figure˜4, highlighting that the effect of block approximation depends strongly on the model and its internal block structure.
+Additional results across other models and datasets are provided in Section˜A.2.1.
+| Identity Translator | Linear Translator | |||||||
+| Approx. | Params. | Accuracy % | GFLOPS | imgs/s | Accuracy % | GFLOPS | imgs/s | |
+| 2 4 | -3.25M | |||||||
+| 9 11 | -3.25M | |||||||
+| 0 1 | -1.62M | |||||||
+| 10 11 | -1.62M | |||||||
+| original | 21.81M | 4.97 | 6349.2 | 4.97 | 6325.6 | |||
+Can entire transformer blocks be approximated without losing accuracy?
+Initial results, reported in Table˜1, support the qualitative analysis and empirically demonstrate that entire vision transformer blocks can be effectively approximated using simpler transformations (e.g., linear projections or, in some cases, the identity function).
+Such approximations reduce both the number of parameters and Giga Floating-Point Operations (GFLOPs), thereby improving throughput (images per second), while incurring only a slight to negligible decrease in downstream task performance.
+For instance, consistent with our earlier analysis, we find that approximating the final block of DEiT-S when using ImageNet1k (e.g., approximating blocks 10 11 or 9 11 with a linear transformation) yields modest performance drops of -0.09% and -3.38%, respectively, while providing substantial efficiency gains.
+Importantly, we also show that even the identity transformation, achieves competitive results, with accuracy drops as small as -0.24% and -5.44%, respectively.
+However, the choice of translator naturally depends on the efficiency-accuracy trade-off: linear translation guarantee in general most reliable accuracy–efficiency balance, whereas the identity yields the leanest training-free approximation when maximum simplicity is required.
+Further methodological details and the full evaluation are presented in Section˜A.1 and Section˜4.2 respectively, while details on the efficiency metrics and additional analysis on those are in Sections˜A.1.5 and A.2.5, respectively.
+4.2 Image Classification Performance
+We evaluate TOAST on image classification tasks using pretrained models of varying sizes (ViT-L, DiNO-B, and DEiT-S) and two benchmark datasets (CIFAR-100F and ImageNet1k).
+Additional results with a broader set of models (ViT-T, ViT-S, ViT-B, ViT-L, DiNO-S, DiNO-B, DEiT-S) and datasets (MNIST, F-MNIST, CIFAR-10, CIFAR-100C) are provided in Section˜A.2.2.
+While in Section˜A.2.6, we complement the quantitative evaluations with qualitative analyses of misclassifications after block approximation, providing further insight into model behavior under TOAST.
+Additional implementation details, including model and dataset specifications, computational resources, and software tools, are provided in Tables˜5 and 6, and Sections˜A.1.4, A.1.5 and A.1.3.
+| Identity | Linear | |||||||
+| Approx. | Params. | Accuracy % | GFLOPs | imgs/s | Accuracy % | GFLOPs | imgs/s | |
+DEiT-S |
+3 4, 9 11 | -4.88M | 3.74 | 7751.4 | 3.80 | 7718.4 | ||
+| 3 4, 9 10 | -3.25M | 4.15 | 7210.9 | 4.21 | 7188.4 | |||
+| 2 3 | -1.62M | 4.56 | 6754.2 | 4.59 | 6736.7 | |||
+| 10 11 | -1.62M | 4.56 | 6752.6 | 4.59 | 6740.5 | |||
+| original | 21.81M | 4.97 | 6349.2 | 4.97 | 6325.6 | |||
+DiNO-B |
+0 4 | -26.00M | 16.32 | 3230.9 | 16.47 | 3227.7 | ||
+| 0 1, 2 3, 4 5 | -19.50M | 18.34 | 2947.0 | 18.80 | 2929.6 | |||
+| 0 1, 2 3 | -13.00M | 20.37 | 2703.9 | 20.67 | 2681.2 | |||
+| 0 1 | -6.50M | 22.39 | 2506.6 | 22.54 | 2487.0 | |||
+| 5 6 | -6.50M | 22.39 | 2503.1 | 22.54 | 2490.6 | |||
+| original | 86.58M | 24.42 | 2321.3 | 24.42 | 2316.5 | |||
+ViT-L |
+2 4, 18 23 | -80.83M | 45.05 | 1654.9 | 45.47 | 1652.8 | ||
+| 17 23 | -69.28M | 47.70 | 1572.4 | 47.90 | 1567.0 | |||
+| 3 4, 19 23 | -57.74M | 50.34 | 1509.9 | 50.75 | 1499.5 | |||
+| 3 4, 20 23 | -46.19M | 52.98 | 1440.4 | 53.39 | 1436.8 | |||
+| 3 4, 21 23 | -34.64M | 55.62 | 1377.2 | 56.03 | 1345.6 | |||
+| 7 8, 15 16 | -23.09M | 58.26 | 1318.2 | 58.67 | 1316.4 | |||
+| 16 17, 22 23 | -23.09M | 58.26 | 1318.8 | 58.67 | 1312.3 | |||
+| 3 4 | -11.55M | 60.90 | 1269.2 | 61.11 | 1270.0 | |||
+| 22 23 | -11.55M | 60.90 | 1267.5 | 61.11 | 1270.9 | |||
+| original | 304.35M | 63.54 | 1219.8 | 63.54 | 1225.2 | |||
+Block approximations in TOAST are calculated via a shared linear, or identity, transformation applied across all tokens and are estimated using a subset of 500 training samples.
+A linear classifier is then trained on top of the frozen backbone with the Adam optimizer (learning rate ), batch size 256, for 5 epochs, over 3 different seeds.
+This setup simulates a realistic scenario where a pretrained feature extractor is adapted to a new dataset unseen during pretraining.
+However, to assess the robustness of our method, we also report the results using the original classification heads (Section˜A.2.4), which confirm the consistency of our findings.
+Are TOAST results competitive?
+As shown in Tables˜3 and 2, TOAST consistently reduces model size and GFLOPs while maintaining, and in some cases improving, image classification accuracy.
+This aligns with our representational analyses in Section˜4.1: for instance, approximating the final block of DEiT-S produces latent representations nearly identical to the original (Figures˜4 and 4), making it an ideal candidate for approximation.
+Even when multiple consecutive blocks are approximated (e.g., 911), models maintain performance comparable to or exceeding the original while significantly reducing parameters.
+This demonstrates that a simple linear transformation, or even the identity in certain cases, is sufficient to capture the functionality of full transformer blocks without additional training, provided the transformation is shared across all tokens.
+Additionally, efficiency gains are notable: throughput (imgs/s) increases while GFLOPs decreases, highlighting practical benefits for deployment, as also shown in Section˜A.2.5.
+Additional results across other models (DiNO-B, ViT-L) and datasets confirm that TOAST generalizes across architectures and scales (Section˜A.2.2).
+Finally, while approximations are easier for simpler datasets (e.g., CIFAR-100F), TOAST still achieves meaningful compression with minimal accuracy loss on complex datasets like ImageNet1k.
+| Identity | Linear | |||||||
+| Approx. | Params. | Accuracy % | GFLOPs | imgs/s | Accuracy % | GFLOPs | imgs/s | |
+DEiT-S |
+3 4, 9 11 | -4.88M | 3.74 | 7755.1 | 3.80 | 7713.7 | ||
+| 9 11 | -3.25M | 4.15 | 7226.6 | 4.18 | 6791.7 | |||
+| 8 9 | -1.62M | 4.56 | 6755.2 | 4.59 | 6739.9 | |||
+| 9 10 | -1.62M | 4.56 | 6692.1 | 4.59 | 6741.3 | |||
+| original | 21.81M | 4.97 | 6349.0 | 4.97 | 6249.4 | |||
+DiNO-B |
+0 4 | -26.00M | 16.32 | 3233.8 | 16.47 | 3204.9 | ||
+| 0 1, 2 3, 4 5 | -19.50M | 18.34 | 2943.1 | 18.80 | 2922.6 | |||
+| 0 1, 2 3 | -13.00M | 20.37 | 2705.6 | 20.67 | 2690.1 | |||
+| 0 1 | -6.50M | 22.39 | 2492.8 | 22.54 | 2493.8 | |||
+| 2 3 | -6.50M | 22.39 | 2484.2 | 22.54 | 2484.6 | |||
+| original | 86.58M | 24.42 | 2315.5 | 24.42 | 2317.3 | |||
+ViT-L |
+2 4, 18 23 | -80.83M | 45.05 | 1655.7 | 45.47 | 1649.6 | ||
+| 17 23 | -69.28M | 47.69 | 1578.8 | 47.90 | 1552.1 | |||
+| 3 4, 19 23 | -57.74M | 50.34 | 1503.6 | 50.75 | 1497.4 | |||
+| 3 4, 20 23 | -46.19M | 52.98 | 1445.2 | 53.39 | 1431.0 | |||
+| 20 23 | -34.64M | 55.62 | 1381.2 | 55.82 | 1372.6 | |||
+| 3 4, 21 23 | -34.64M | 55.62 | 1376.7 | 56.03 | 1372.7 | |||
+| 20 22 | -23.09M | 58.26 | 1317.5 | 58.47 | 1314.6 | |||
+| 3 4, 21 22 | -23.09M | 58.26 | 1315.8 | 58.67 | 1317.6 | |||
+| 20 21 | -11.55M | 60.90 | 1268.5 | 61.11 | 1266.7 | |||
+| 21 22 | -11.55M | 60.90 | 1270.7 | 61.11 | 1269.2 | |||
+| original | 304.35M | 63.54 | 1223.1 | 63.54 | 1224.3 | |||
+Additional results across models and datasets are provided in Tables˜7, 8, 9 and 10.
+To assess scalability, we applied TOAST to ViT-L.
+Approximating selected blocks, e.g., 17 23, reduces the parameter count by 69.3M, lowers GFLOPs from 63.54 to 47.79, and increases throughput from 1223.1 to 1578.8 imgs/s, while incurring a minimal accuracy drop of 1.16%.
+This shows TOAST’s utility in balancing substantial computational savings with a modest performance trade-off, even in large models.
+Are 500 training samples enough?
+We study the sensitivity of block approximation to the number of training samples using DiNO-B and DEiT-S on ImageNet1k.
+As shown in Figure˜5, performance typically plateaus quickly: 500 samples are sufficient to obtain stable and reliable approximations.
+Increasing the sample count beyond this threshold provides only marginal gains, while substantially fewer samples lead to noticeable degradation.
+Interestingly, when the representational spaces of consecutive blocks are already highly aligned, even as few as 10 or 50 samples suffice to achieve competitive approximations.
+Conversely, for blocks that are harder to approximate, such as the early layers of DEiT-S (e.g., 01), even 4000 samples are insufficient to estimate a linear transformation that maintains competitive performance.
+We highlight that these results are obtained on ImageNet1k, which contains 1000 classes.
+The 500 samples represent only a small subset of the class space, yet reliable approximations are still achieved.
+This indicates that TOAST primarily captures the block-level structure of representations rather than requiring exhaustive coverage of all classes.
+Consequently, TOAST could be practical also in scenarios where a large labeled datasets is limited.
+What if a more complex transformation is used?
+We evaluate whether deeper approximators improve downstream task performance.
+Specifically, we compare TOAST (Identity and Linear) to MultiLayer Perceptron (MLP) and Residual MLP, trained for 300 steps with Adam (learning rate ).
+These more complex transformation, as for Identity and Linear, are applied across all tokens, and estimated using a subset of 500 training samples.
+Results in Table˜4 show a consistent trend for ViT-L on both ImageNet1k and CIFAR-100F: the linear transformation provides the most reliable trade-off across datasets.
+On CIFAR-100F, linear often achieves the best or near-best accuracy (e.g., 2122: vs.
+for Res-MLP and for MLP), while remaining training-free, thus more efficient.
+On ImageNet1k, the gap becomes even clearer: for the same blocks linear reaches , while Res-MLP and MLP reach and , respectively.
+Additionally, also Linear obtain competitive results.
+TOAST operates in closed form, requires no optimization, and consistently achieves strong efficiency–accuracy trade-offs.
+These findings confirm that a simple linear transformation is sufficient to approximate transformer blocks in most settings, with deeper translators offering little benefit despite their higher cost.
+| Accuracy | ||||||
+| Approx. | Params. | Identity | Linear | MLP | Res-MLP | |
+CIFAR-100F |
+3 4, 20 23 | -46.19M | ||||
+| 17 23 | -69.28M | |||||
+| 3 4, 19 23 | -57.74M | |||||
+| 20 23 | -34.64M | |||||
+| 20 22 | -23.09M | |||||
+| 3 4, 21 22 | -23.09M | |||||
+| 20 21 | -11.55M | |||||
+| 21 22 | -11.55M | |||||
+| original | 304.35M | |||||
+ImageNet1k |
+3 4, 20 23 | -46.19M | ||||
+| 17 23 | -69.28M | |||||
+| 3 4, 19 23 | -57.74M | |||||
+| 20 23 | -34.64M | |||||
+| 20 22 | -23.09M | |||||
+| 3 4, 21 22 | -23.09M | |||||
+| 20 21 | -11.55M | |||||
+| 21 22 | -11.55M | |||||
+| original | 304.35M | |||||
 5 Limitations and future work
-While TBA demonstrates significant promise in efficiently approximating transformer blocks, our current investigation has primarily focused on vision transformer architectures and their application to classification tasks.
-Future research will explore the applicability of TBA to other modalities (e.g., text) and to diverse downstream tasks (e.g., image reconstruction).
-Such an expansion will be crucial for testing the universality of the observed block-similarity phenomena and assessing TBA’s adaptability.
+While TOAST efficiently approximates transformer blocks, our current investigation has primarily focused on vision transformer architectures and their application to classification tasks.
+Future research will explore the applicability of TOAST to other modalities (e.g., text) and to diverse downstream tasks (e.g., image reconstruction).
+Such an expansion will be crucial for testing the universality of the observed block-similarity phenomena and assessing TOAST’s adaptability.
 Furthermore, we aim to expand the analysis of these block-level similarities.
-This involves investigating redundancies at finer granularities, such as within individual attention heads or feed-forward network layers, and potentially at coarser granularities across larger segments of the network.
-Another important direction is to examine more in detail how these similarity patterns arise depending on token type (e.g., CLS token versus last token).
+This involves investigating redundancies at finer granularities, such as within individual attention heads or feed-forward layers, and consistently and automatically identifying which blocks to approximate, resulting in minimal degradation of downstream task performance.
+Another important direction is to examine more in detail how these similarity patterns arise depending on token type (e.g., CLS token versus mean token).
 The analysis could lead to even more refined and context-aware approximation strategies, further enhancing model efficiency.
 6 Conclusion
-In this work, we first analyze the emergence of consistent block-wise representation similarities within pretrained foundation models and then propose a method to leverage these similarities to obtain smaller yet performant models.
-To this end, we propose Transformer Blocks Approximation (TBA), a novel method for identifying and efficiently approximating similar transformer blocks using a simple linear transformation, without requiring additional training or fine-tuning.
-Our extensive empirical evaluations across multiple pretrained vision models and datasets validate that TBA significantly reduces model parameters while maintaining, and sometimes even improving, downstream task performance.
-Notably, the learned linear transformations exhibit strong generalization capabilities across datasets, enabling lightweight model adaptation.
-Furthermore, TBA’s straightforward linear approach often achieves better results than existing strategies like block skipping, and can be more effective than complex, trained approximations.
-TBA thus offers a practical and efficient method for streamlining foundation models, making them more computationally accessible.
+In this work, we first analyze the emergence of consistent block-wise representation similarities within pretrained foundation models and then propose a method to leverage these similarities to obtain smaller and more efficient yet performant models.
+To this end, we propose Transformer Optimization using Adaptive and Simple Transformations (TOAST), a novel method for easily approximate entire transformer blocks using a simple transformation, without requiring additional training or fine-tuning.
+Our extensive empirical evaluations across multiple pretrained vision models and datasets validate that TOAST significantly reduces model parameters while maintaining, and sometimes even improving, downstream task performance.
+Furthermore, TOAST’s straightforward linear approach often achieves better results than existing strategies like block skipping, and can be more effective than complex, trained approximations.
+TOAST thus offers a practical and efficient method for streamlining foundation models, making them more computationally accessible, and towards deployment in resource-constrained scenarios such as on-device settings.
 Acknowledgments
@@
 A.1 Implementation details
-This section details the experiments conducted in Section 4, providing information to reproduce them. Additionally, we provide the code as a zip file in the supplementary material.
+This section details the experiments conducted in Section˜4, providing information to reproduce them. Additionally, we provide the code as a zip file in the supplementary material.
 A.1.1 Models and Datasets
-Table 6 contains the full list of the pretrained models, while Table 7 contains dataset information.
+Table˜5 contains the full list of the pretrained models, while Table˜6 contains dataset information.
 | Modality | HuggingFace Model Name | Alias | Enc. Dim |
@@
 DVC (Kuprieiev et al., 2022), for data versioning;
+fvcore analysis library , for calculating GFLOPs;
 A.1.4 Computational resources
@@
 All the other experiments utilized an NVIDIA GeForce RTX 5090 GPU with 31 GB of memory.
+A.1.5 Efficiency metrics
+We evaluated model efficiency using two primary metrics.
+GFLOPs were used to measure the hardware-independent theoretical complexity of a single forward pass, calculated using the fvcore analysis library.
+Throughput, measured in samples per second, was used to quantify the practical, hardware-dependent inference speed.
+This was benchmarked by averaging the wall-clock time over numerous iterations on a single NVIDIA H100 GPU with a consistent batch size of 256.
 A.2 Additional Experiments
-This section presents supplementary experiments to extend those detailed in Section 4.
-A.2.1 Block Similarities
-The analysis in Section 4.1 reveals distinct block-wise similarity patterns within pretrained foundation models.
-The analysis utilizes the mean of token representations within each block to compute similarities.
-However, as detailed in Figure 5, these characteristic similarity patterns persist and are qualitatively very similar when using only the CLS token’s representation from each block.
-This suggests a degree of interchangeability for this type of analysis.
-Finally, to assess the impact of the similarity measure itself, we conducted an ablation study correlating various metrics with downstream task performance on ViT-S and ImageNet1k (Figure 6).
-Our findings reveal that even simple metrics, such as MSE or cosine similarity, when used to identify block-level similarities, exhibit a strong correlation with final accuracy.
-These results underscore the robustness and flexibility of leveraging such block similarities, indicating compatibility with a range of similarity or distance functions capable of capturing these inter-block relationships.
-| Approx. | Params. | ImageNet1k |
-| 1 5 | 15.31M | |
-| 2 5 | 16.94M | |
-| 7 10 | 16.94M | |
-| 1 3 | 18.56M | |
-| 3 5 | 18.56M | |
-| 2 4 | 18.56M | |
-| 9 11 | 18.56M | |
-| 2 3 | 20.19M | |
-| 3 4 | 20.19M | |
-| 4 5 | 20.19M | |
-| 9 10 | 20.19M | |
-| - | 21.82M |
-A.2.2 PCA Visualization
-In Section 4.1 we show the effect of TBA on latent representations.
-In this section, we provide additional visualization using PCA for DiNO-S, DEiT-S, ViT-S, with different datasets and approximating both early and late blocks (see Figures 10, 11, 9, 7 and 8).
-A.2.3 Image Classification
-This section presents additional experiments that complement and extend those detailed in Section 4.2. Datasets and models are the ones detailed in Tables 7 and 6.
-| Approx. | Params. | MNIST | F-MNIST | CIFAR-10 | CIFAR-100C | CIFAR-100F | ImageNet1k |
-| 1 5 | 15.31M | ||||||
-| 2 5 | 16.94M | ||||||
-| 7 10 | 16.94M | ||||||
-| 1 3 | 18.56M | ||||||
-| 2 4 | 18.56M | ||||||
-| 3 5 | 18.56M | ||||||
-| 1 2, 3 4 | 18.56M | ||||||
-| 1 2, 4 5 | 18.56M | ||||||
-| 0 1 | 20.43M | ||||||
-| 1 2 | 20.43M | ||||||
-| 2 3 | 20.43M | ||||||
-| 3 4 | 20.43M | ||||||
-| 4 5 | 20.43M | ||||||
-| 5 6 | 20.43M | ||||||
-| 6 7 | 20.43M | ||||||
-| 7 8 | 20.43M | ||||||
-| 8 9 | 20.43M | ||||||
-| 9 10 | 20.43M | ||||||
-| 10 11 | 20.43M | ||||||
-| original | 22.06M |
+This section presents supplementary experiments to extend those detailed in Section˜4.
+A.2.1 Latent Analysis
+This section extend the analysis conducted in Section˜4.1, to analyze block-wise internal similarities, to additional models of different dimensionality: ViT-T, ViT-S, ViT-B and DiNO-S.
+Additionally, we provide visualization using PCA for DiNO-S, DEiT-S, ViT-S, with different datasets and approximating both early and late blocks (see Figures˜10, 11, 9, 7 and 8).
+A.2.2 Image Classification
+This section presents additional experiments that complement and extend those detailed in Section˜4.2. Datasets and models are the ones detailed in Tables˜6 and 5.
 | Approx. | Params. | MNIST | F-MNIST | CIFAR-10 | CIFAR-100C | CIFAR-100F | ImageNet1k |
@@
 | - | 86.39M | |||||
-| Accuracy | |||||||
-| Approx. | Params. | MNIST | F-MNIST | CIFAR-10 | CIFAR-100C | CIFAR-100F | ImageNet1k |
-| 112 | -127.02M | ||||||
-| 04 | -46.19M | ||||||
-| 15 | -46.19M | ||||||
-| 25 | -34.64M | ||||||
-| 710 | -34.64M | ||||||
-| 02, 34 | -34.64M | ||||||
-| 12, 34 | -23.09M | ||||||
-| 13 | -23.09M | ||||||
-| 24 | -23.09M | ||||||
-| 35 | -23.09M | ||||||
-| 57 | -23.09M | ||||||
-| 01 | -11.55M | ||||||
-| 12 | -11.55M | ||||||
-| 23 | -11.55M | ||||||
-| 34 | -11.55M | ||||||
-| 45 | -11.55M | ||||||
-| 56 | -11.55M | ||||||
-| 67 | -11.55M | ||||||
-| 78 | -11.55M | ||||||
-| 89 | -11.55M | ||||||
-| 910 | -11.55M | ||||||
-| 1011 | -11.55M | ||||||
-| 1112 | -11.55M | ||||||
-| 1213 | -11.55M | ||||||
-| 1314 | -11.55M | ||||||
-| 1415 | -11.55M | ||||||
-| 1516 | -11.55M | ||||||
-| 1617 | -11.55M | ||||||
-| 1718 | -11.55M | ||||||
-| 1819 | -11.55M | ||||||
-| 1920 | -11.55M | ||||||
-| 2021 | -11.55M | ||||||
-| 2122 | -11.55M | ||||||
-| 2223 | -11.55M | ||||||
-| original | 304.35M | ||||||
-A.2.4 Transformation Generalization
-In this section, we extend the analysis to evaluate the model’s ability to generalize transformations across datasets to different approximation blocks and models (i.e., DiNO-S).
-| Accuracy | ||||||
-| Encoder | Approx. | Source | MNIST | CIFAR-10 | CIFAR-100C | CIFAR-100F |
-| ViT-S | 2 3 | MNIST | ||||
-| CIFAR-10 | ||||||
-| CIFAR-100 | ||||||
-| 3 4 | MNIST | |||||
-| CIFAR-10 | ||||||
-| CIFAR-100 | ||||||
-| 4 5 | MNIST | |||||
-| CIFAR-10 | ||||||
-| CIFAR-100 | ||||||
-| 9 10 | MNIST | |||||
-| CIFAR-10 | ||||||
-| CIFAR-100 | ||||||
-| 1 3 | MNIST | |||||
-| CIFAR-10 | ||||||
-| CIFAR-100 | ||||||
-| 3 5 | MNIST | |||||
-| CIFAR-10 | ||||||
-| CIFAR-100 | ||||||
-| 2 4 | MNIST | |||||
-| CIFAR-10 | ||||||
-| CIFAR-100 | ||||||
-| 8 10 | MNIST | |||||
-| CIFAR-10 | ||||||
-| CIFAR-100 | ||||||
-| 9 11 | MNIST | |||||
-| CIFAR-10 | ||||||
-| CIFAR-100 | ||||||
-| 2 5 | MNIST | |||||
-| CIFAR-10 | ||||||
-| CIFAR-100 | ||||||
-| 7 10 | MNIST | |||||
-| CIFAR-10 | ||||||
-| CIFAR-100 | ||||||
-| 1 5 | MNIST | |||||
-| CIFAR-10 | ||||||
-| CIFAR-100 | ||||||
-| Accuracy | ||||||
-| Encoder | Approx. | Source | MNIST | CIFAR-10 | CIFAR-100C | CIFAR-100F |
-| DiNO-S | 2 3 | MNIST | ||||
-| CIFAR-10 | ||||||
-| CIFAR-100 | ||||||
-| 3 4 | MNIST | |||||
-| CIFAR-10 | ||||||
-| CIFAR-100 | ||||||
-| 4 5 | MNIST | |||||
-| CIFAR-10 | ||||||
-| CIFAR-100 | ||||||
-| 9 10 | MNIST | |||||
-| CIFAR-10 | ||||||
-| CIFAR-100 | ||||||
-| 1 3 | MNIST | |||||
-| CIFAR-10 | ||||||
-| CIFAR-100 | ||||||
-| 3 5 | MNIST | |||||
-| CIFAR-10 | ||||||
-| CIFAR-100 | ||||||
-| 2 4 | MNIST | |||||
-| CIFAR-10 | ||||||
-| CIFAR-100 | ||||||
-| 8 10 | MNIST | |||||
-| CIFAR-10 | ||||||
-| CIFAR-100 | ||||||
-| 9 11 | MNIST | |||||
-| CIFAR-10 | ||||||
-| CIFAR-100 | ||||||
-| 2 5 | MNIST | |||||
-| CIFAR-10 | ||||||
-| CIFAR-100 | ||||||
-| 7 10 | MNIST | |||||
-| CIFAR-10 | ||||||
-| CIFAR-100 | ||||||
-| 1 5 | MNIST | |||||
-| CIFAR-10 | ||||||
-| CIFAR-100 | ||||||
-A.2.5 Comparison
-In this section, we provide additional results for the SkipAt baseline.
-| Accuracy | |||||
-| Skip | MNIST | F-MNIST | CIFAR-10 | CIFAR-100C | CIFAR-100F |
-| 1 5 | |||||
-| 2 5 | |||||
-| 7 10 | |||||
-| 1 3 | |||||
-| 3 5 | |||||
-| 2 4 | |||||
-| 8 10 | |||||
-| 9 11 | |||||
-| 0 1 | |||||
-| 1 2 | |||||
-| 2 3 | |||||
-| 3 4 | |||||
-| 4 5 | |||||
-| 5 6 | |||||
-| 6 7 | |||||
-| 7 8 | |||||
-| 8 9 | |||||
-| 9 10 | |||||
-| 10 11 | |||||
-| original | |||||
+A.2.3 Zero-shot Image Classification
+To further assess the effectiveness of our approach, we evaluate TOAST in a zero-shot image classification setting.
+This evaluation utilizes the OpenCLIP-ViT-B model (Radford et al., 2021), which was pretrained on LAION-2B Schuhmann et al.
+(2022), with ImageNet1k serving as the downstream evaluation dataset.
+| Params. | Approx. | Accuracy | |
+| -6.49M | 0 1 | 57.93 | - |
+| 1 2 | 64.20 | - | |
+| 2 3 | 66.35 | -5.51% | |
+| 3 4 | 64.65 | - | |
+| 4 5 | 64.86 | - | |
+| 5 6 | 58.05 | - | |
+| 6 7 | 61.56 | - | |
+| 7 8 | 58.53 | - | |
+| 8 9 | 52.32 | - | |
+| 9 10 | 59.21 | - | |
+| 10 11 | 22.64 | - | |
+| 149.07M | original | 70.21 | – |
+The analysis is conducted only on the base version, as larger versions (e.g., OpenCLIP-ViT-L or OpenCLIP-ViT-H) contain too many parameters and are thus beyond the scope of this paper.
+As in previous experiments, the model remains frozen, and block approximations are computed using a shared linear transformation applied across all tokens, based on a subset of 3,000 training samples.
+Importantly, we apply these approximations only to the vision encoder, leaving the text encoder unchanged.
+We follow the standard ImageNet1k prompt templates.
+The results in Table˜11 lead to the conclusion that the impact on zero-shot accuracy is highly dependent on the targeted block’s position.
+The choice of which blocks to approximate is therefore crucial.
+For instance, approximating an early block (e.g., 1 2 or 2 3) results in a modest accuracy drop (e.g., 5.51%), yielding a competitive model with fewer parameters.
+In contrast, approximating the final block (i.e., 10 11) causes a catastrophic performance collapse of 67.75%.
+Meaning that, for OpenCLIP-ViT-B, later layers in the vision encoder appear to capture uniquely critical information for zero-shot generalization that cannot be effectively replicated by earlier ones.
+To the best of our knowledge, this work is the first to investigate training-free model size reduction in this challenging setting.
+A.2.4 Evaluation with original classification heads
+| Encoder | Approximation | Accuracy |
+DEiT-S |
+78.87 | |
+| 78.72 | ||
+| 78.39 | ||
+| 75.29 | ||
+| 74.36 | ||
+| 72.34 | ||
+| 69.30 | ||
+| 66.12 | ||
+| original | 79.66 | |
+ViT-S |
+78.25 | |
+| 74.50 | ||
+| 71.23 | ||
+| 70.46 | ||
+| 66.18 | ||
+| 59.69 | ||
+| 58.36 | ||
+| 23.85 | ||
+| original | 79.86 |
+As mentioned in the main paper, our primary evaluation involves training a new linear classifier on top of the frozen model backbone to simulate a realistic transfer learning scenario.
+However, the original papers for DEiT-S Touvron et al.
+(2021) and ViT-S Beyer et al.
+(2022) report performance using the classification head that was part of the original pre-training.
+To confirm that our conclusions are robust and not an artifact of our evaluation protocol, we conducted an additional set of experiments using the official, pre-trained classification heads from the original model checkpoints.
+In this setup, we do not train a new classifier; we simply evaluate the accuracy of the frozen, approximated models using their original heads.
+The results, presented in Table˜12, are fully consistent with the main conclusions of our paper.
+They confirm that our block approximation method provides a favorable accuracy-efficiency trade-off, even when evaluated with the original model heads.
+The relative drop in accuracy when approximating different layers follows the same patterns observed in our primary experiments, reinforcing the validity of our approach.
+A.2.5 Computational efficiency vs. accuracy
+To quantify the effectiveness of different approximation methods, we analyze the trade-off between downstream accuracy and computational cost.
+Figure˜12 presents this analysis on a DiNO-B model using both CIFAR-100F and ImageNet1k against three standard efficiency metrics: parameter count, GFLOPs, and inference throughput.
+Across all metrics, the proposed linear translator (green) establishes a more favorable Pareto frontier compared to the baseline identity-based approach (blue).
+This indicates that for any given efficiency budget (e.g., a specific GFLOPs target), the linear translator consistently yields a model with higher accuracy.
 A.2.6 Analysis of misclassifications
 In this section, we examine changes in per-class accuracy and misclassification patterns.
-As shown in Figure 12, models behave differently at block approximations.
+As shown in Figure˜13, models behave differently at block approximations.
 DiNO-S remains remarkably stable across blocks and classes, with the only degradation appearing for classes dog (when approximating blocks 10 or 11) and deer (for block 10 approximation).
@@
 Instead, the most noticeable hit occurs for class cat when the earlier blocks are approximated.
-For DEiT-S, several mid-to-late block approximations improve accuracy for various classes, whereas the very first block causes a clear relative decline in nearly every class.
-These observations suggest strategies like preferring late-block approximation for DEiT-S, or reserving extra samples for the linear transformation in order to recover the accuracy of difficult classes for the model.
+For DEiT-S, several mid‑to‑late block approximations improve accuracy for various classes, whereas the very first block causes a clear relative decline in nearly every class.
+These observations suggest strategies like preferring late‑block approximation for DEiT-S, or reserving extra samples for the linear transformation in order to recover the accuracy of difficult classes for the model.
 In order to further investigate how the predictions change while approximating blocks, we plot the difference in the normalized confusion matrix before and after the approximation.
-In Figure 13, we show the delta confusion matrix for DEiT-S on CIFAR-100C.
+In Figure˜14, we show the delta confusion matrix for DEiT-S on CIFAR-100C.
 Also, here we can see how approximating the very first block makes the model puzzling and lose per-class accuracy (i.e., negative delta along the diagonal).
 On the other hand, approximating the last block acts as a regularizer, resulting in an overall gain in the per-class accuracy and, as a consequence, fewer misclassifications (negative deltas off-diagonal).
-This supports results shown in Figures 12 and 1.
-Additionally, Figure 14 shows representative CIFAR-10 images that become misclassified after approximating a block of ViT-S.
-The patterns we observe mirror the trends in Figures 12 and 13: when approximating earlier blocks, we observe many images belonging to class cat to be misclassified.
+This supports results shown in Figures˜13 and 3.
+Additionally, Figure˜15 shows representative CIFAR-10 images that become misclassified after approximating a block of ViT-S.
+The patterns we observe mirror the trends in Figures˜13 and 14: when approximating earlier blocks, we observe many images belonging to class cat to be misclassified.
 Instead, when approximating later blocks, we observe images of the class dog to be misclassified.